### 31 Mar 2026 v1

Abstract
Next-generation radio astronomy surveys are delivering millions of resolved sources, but scientific exploitation is increasingly limited by robust, scalable morphology analysis across heterogeneous telescopes and imaging pipelines. We present STRADAViT, a self-supervised Vision Transformer (ViT) continued-pretraining framework for learning transferable encoders from radio astronomy imagery. The framework combines mixed-survey data curation, radio astronomy-aware training-view generation, and a ViT-MAE backbone with optional register tokens. It supports reconstruction-only, contrastive-only, and two-stage branches. Our pretraining dataset comprises $512\times 512$ radio astronomy images drawn from four complementary sources (MeerKAT, ASKAP, LOFAR/LoTSS, and SKA). We evaluate transfer with linear probing and fine-tuning on three morphology benchmarks spanning binary and multi-class settings (MiraBest, LoTSS DR2, and Radio Galaxy Zoo). Relative to the ViT-MAE initialization used for continued pretraining, the best two-stage models improve Macro-F1 in all reported linear-probe settings and in two of three fine-tuning settings, with the largest gain on RGZ DR1. Relative to DINOv2, gains are selective rather than universal: the best two-stage models surpass the strongest DINOv2 baseline on LoTSS DR2 and RGZ DR1 under linear probing, and on MiraBest and RGZ DR1 under fine-tuning. A targeted DINOv2-initialized HCL ablation further indicates that the adaptation recipe is not specific to the ViT-MAE starting point. The ViT-MAE-based STRADAViT checkpoint nevertheless remains the preferred release because it combines competitive transfer with substantially lower token count and downstream cost than the DINOv2-based alternative. These results indicate that radio astronomy-aware view generation and staged continued pretraining can provide a stronger domain-adapted starting point than out-of-the-box ViT checkpoints for radio astronomy transfer, especially when representation quality is assessed through linear probing.
keywords:
[…]
{highlights}
Mixed-survey self-supervised dataset spanning MeerKAT, ASKAP, LoTSS, and SKA-derived cutouts
Radio astronomy-aware view generation for sparse and heterogeneous radio astronomy images
Controlled comparison of reconstruction-only, contrastive-only, and two-stage continued pretraining
Two-stage STRADAViT improves on the ViT-MAE starting point in all reported linear-probe settings and in most fine-tuning settings
Domain-adapted transfer is strongest under frozen-backbone evaluation, with selective gains against DINOv2 baselines
1 Introduction
Radio astronomy is entering an era of high-volume imaging surveys, with instruments such as MeerKAT, ASKAP and LOFAR producing increasingly deep and wide observations, and future facilities such as the SKA expected to extend both scale and sensitivity. These surveys contain diverse morphologies and imaging artefacts that can masquerade as morphology. The scientific return is often constrained by the availability of reliable, scalable image-based analysis tools that generalize across instruments and data products, rather than by raw data volume.
In radio astronomy, supervised deep learning faces two recurring constraints. First, high-quality labels for morphology are expensive and typically derived from expert visual inspection or citizen science. Second, models trained on one survey or imaging pipeline can degrade when applied to others due to differences in angular resolution, $uv$-coverage, deconvolution, noise statistics, dynamic range, and calibration/imaging conventions. This motivates approaches that exploit abundant unlabeled radio astronomy data while explicitly targeting cross-telescope robustness.
In mainstream computer vision, transferable representations are increasingly learned via self-supervised pretraining at scale and then adapted to diverse downstream tasks. Vision Transformers (ViTs) build on the Transformer paradigm (Vaswani et al., 2017) and have proven effective when scaled and pretrained appropriately (Dosovitskiy et al., 2021). Self-supervised objectives such as masked image modeling (He et al., 2022) and contrastive or invariance-based learning (Chen et al., 2020; He et al., 2020; Grill et al., 2020; Bardes et al., 2022) enable representation learning without labels, while recent work demonstrates strong ViT features learned purely from self-supervision (Caron et al., 2021; Oquab et al., 2023). A key open question for radio astronomy is how to adapt these pretraining ideas under the constraints of scientific imaging: single-channel data, survey-dependent intensity distributions, and heterogeneous instrumental/imaging systematics.
This paper presents STRADAViT, developed as part of STRADA (Self-supervised Transformers for Radio Astronomy Discovery Algorithms). The framework is organized around three components: (i) a mixed-survey self-supervised pretraining dataset drawn from four complementary radio astronomy image sources, (ii) radio astronomy-aware training-view generation for masked reconstruction and contrastive learning, and (iii) a continued-pretraining framework built on ViT-MAE. The framework supports reconstruction-only, contrastive-only, and two-stage branches.
Our primary contributions are:
  • a mixed-survey $512\times 512$ radio astronomy cutout dataset spanning four complementary sources;
  • radio astronomy-aware training-view generation that anchors self-supervised views to informative source regions instead of relying on naive random crops;
  • a ViT-MAE-based self-supervised framework with controlled ablations of reconstruction losses, hard-negative-aware contrastive losses, and register tokens, evaluated through reconstruction-only, contrastive-only, and two-stage continued pretraining;
  • a reproducible transfer protocol (linear probing and fine-tuning) across three public morphology benchmarks, with analysis emphasizing the probe–fine-tune gap and dataset sensitivity.
Results show that radio astronomy-aware views and staged continued pretraining improve a ViT-MAE starting point and yield more transferable representations for sparse radio astronomy data, although strong off-the-shelf vision baselines remain competitive.
Section 2 reviews the relevant vision and radio-astronomy literature. Sections 3 and 4 describe the pretraining and evaluation datasets. Sections 57 define the model, training branches, and pretraining objectives. Transfer results are then presented under linear probing and fine-tuning. We then present classwise confusion-pattern comparisons against the starting-point baselines, followed by an overall interpretation of the branch comparison and the conclusion.
2 Literature Survey
This survey covers the transferable-representation literature most directly relevant to the present design choices.
2.1 Backbones for Transferable Visual Representations
Convolutional neural networks remain strong imaging baselines, particularly in residual form (He et al., 2016), but Transformer-based backbones now dominate large-scale representation learning (Vaswani et al., 2017; Dosovitskiy et al., 2021). In specialized imaging domains, backbone choice interacts closely with pretraining objective, dataset composition, and normalization.
2.2 Self-Supervised Learning: Contrastive Invariances and Masked Modeling
Self-supervised learning replaces manual labels with surrogate objectives. Contrastive methods such as SimCLR and MoCo learn invariances across augmented views (Chen et al., 2020; He et al., 2020), whereas related methods modify the handling of negatives or regularize feature statistics (Grill et al., 2020; Bardes et al., 2022). Masked image modeling provides a complementary signal: MAE reconstructs masked content from partial observations and is effective when image statistics differ substantially from those of natural images (He et al., 2022). DINO-style methods and DINOv2 further show that training recipe and architectural details, including register tokens, materially affect transfer (Caron et al., 2021; Oquab et al., 2023). These two objective families motivate the staged design studied here.
2.3 From General-Purpose Pretraining to Domain-Specific Transfer
General-purpose pretraining can yield strong transferable features (Radford et al., 2021), but radio astronomy differs markedly from natural-image domains: images are typically single-channel, have survey-dependent noise and dynamic range, and contain instrument- and pipeline-specific artefacts. Effective transfer therefore depends on representative unlabeled data, domain-appropriate normalization, and augmentations that preserve morphology rather than photographic appearance.
2.4 Radio Astronomy: From Source Finding to Transferable Encoders
Beyond morphology classification, radio astronomy has a long tradition of source finding and characterization under survey-specific artefacts and systematics, as illustrated by CAESAR and related practical pipelines (Riggi et al., 2016, 2019). Recent work extends this toward broad reviews of deep learning for radio classification (Riggi et al., 2024), self-supervised transfer and benchmark studies (Riggi et al., 2024; Cecconello et al., 2025), and multimodal systems (Riggi et al., 2025; Drozdova et al., 2025). Additional SSL studies now cover source classification, FR I/FR II transfer, foundation-style pretraining on Radio Galaxy Zoo, and self-supervised learning on MeerKAT continuum images (Baron Pérez et al., 2025; Buatthaisong et al., 2025; Slijepcevic et al., 2023; Lastufka et al., 2024).
2.5 Radio Morphology Datasets and Benchmark Labels
Radio morphology benchmarks remain anchored in expert or citizen-science labels. LoTSS DR2 provides broad morphology labels and feature flags for bright, extended sources (Horton et al., 2025); MiraBest provides curated FR I/FR II splits (Porter and Scaife, 2023); and Radio Galaxy Zoo operationalizes large-scale citizen-science labeling for more complex multi-class settings (Wong et al., 2024). Across these datasets, labels cover only part of the available data and label spaces are not harmonized across surveys, which makes cross-survey evaluation essential.
2.6 Positioning Relative to Recent Radio Astronomy SSL Literature
Relative to recent radio astronomy SSL studies (Riggi et al., 2024; Cecconello et al., 2025; Lastufka et al., 2024; Baron Pérez et al., 2025; Buatthaisong et al., 2025; Slijepcevic et al., 2023), STRADAViT emphasizes four points: mixed-survey pretraining, ROI-aware view generation for sparse cutouts, a staged reconstruction-to-contrastive objective, and a controlled ViT-only architectural scope. The aim is cross-survey reuse across multiple downstream label spaces rather than optimization for a single benchmark.
3 Unsupervised Learning Dataset
Our self-supervised pretraining dataset is constructed to reflect the heterogeneity of contemporary radio astronomy imaging across instruments and imaging pipelines. We assemble $512\times 512$ continuum-image cutouts from four complementary sources: MeerKAT MGCLS DR1, ASKAP continuum images accessed via the CSIRO ASKAP Science Data Archive, LOFAR LoTSS continuum mosaics, and simulated continuum images from the first SKA Science Data Challenge (SDC1) DR1 (Bonaldi et al., 2020).
3.1 Cutout Extraction and Basic Filtering
All inputs are treated as single-channel images stored in FITS. To obtain a uniform pixel-space resolution, we tile each survey image into non-overlapping $512\times 512$ cutouts (stride equal to tile size), discarding partial tiles at image edges. When the input product is a cube, we iterate over its leading axes and extract cutouts from each 2D plane.
We apply basic filtering to exclude uninformative or corrupted data from the pretraining set. Cutouts containing NaN/Inf values or consisting entirely of zeros are discarded. We additionally skip malformed FITS products (e.g., missing image arrays, empty arrays, or non-numeric pixel types).
3.2 Cutout Normalization
Radio cutouts can vary substantially in absolute intensity scale and dynamic range across surveys. Since our downstream focus is morphology-centric, we apply a standard astronomy contrast normalization to each cutout independently using Astropy’s ZScaleInterval (IRAF-style ZScale). For each image, ZScale estimates display-style lower/upper limits and returns a linearly scaled image; we then clip the result to $[0,1]$. Any non-finite values are set to zero, and pathological constant cutouts are mapped to zeros. This preprocessing does not preserve absolute flux calibration by design. For compatibility with standard vision backbones, the resulting single-channel image is replicated to three channels after scaling.
3.3 Dataset Composition
Table 1 summarizes the resulting dataset size after preprocessing and filtering.
Table 1: Composition of the self-supervised pretraining dataset (all images are $512\times 512$ cutouts). “Ratio” denotes the fraction of total cutouts contributed by each source.
Source Cutouts Ratio
MGCLS (MeerKAT), DR1 139500 0.24
ASKAP (ASDA) 261390 0.44
LoTSS (LOFAR) 177476 0.30
SKA SDC1, DR1 12288 0.02
Total 590654 1.00
Figure 1: Reconstruction-branch masked-reconstruction views. The first column shows the standardized parent cutout (after the per-image ZScale contrast stretch described in Section 3). Subsequent columns show example ROI-aligned crops produced by our single-view strategy (Section 6); with probability $p_{\mathrm{global}}=0.2$ the strategy instead samples a wide-field crop from the full cutout. The selected view is then transformed by progressively applied, morphology-preserving augmentations.
Figure 2: Contrastive-branch multi-view examples produced by our on-the-fly augmenter. The first column shows the standardized parent cutout. Each subsequent pair of columns illustrates two correlated views sampled from the same cutout and anchored to the same object-centric ROI (Section 7): a wider/global view with mild, morphology-preserving augmentations and an additional view with stronger corruptions. For training we sample exactly $V=2$ views per cutout per forward pass, so each step uses one such pair per cutout; the multiple pairs shown here are independent draws for visualization. In the contrastive loss, the two views of the same cutout form the positives, while all other views in the (distributed) micro-batch act as negatives.
3.4 Online View Generation and Extensibility
All views used for self-supervision are generated on-the-fly during training (Sections 6 and 7), rather than being precomputed. As a consequence, a given cutout is typically encountered under many different ROI selections, crop scales, and augmentations over the course of training, which increases effective data diversity without requiring additional offline curation. This runtime design is also intended to make the pretraining dataset extensible: additional telescope or simulation sources can be incorporated by adding new mosaics (or cutouts) that satisfy the same basic FITS tiling and preprocessing interface, without redesigning survey-specific curation rules.
4 Evaluation Datasets
We evaluate transfer learning on three public, morphology-focused radio astronomy imaging datasets that span different telescopes, labeling paradigms, and class granularities: MiraBest (Porter and Scaife, 2023), the LoTSS DR2 visual-classification sample of Horton et al. (2025), and Radio Galaxy Zoo DR1 (Wong et al., 2024). Together, these datasets allow us to probe binary and multi-class settings and to test robustness across heterogeneous imaging products.
Table 2 summarizes the number of retained samples in each evaluation dataset after applying the dataset selection and filtering criteria described below.
Table 2: Evaluation datasets and retained sample counts.
Dataset Samples
Radio Galaxy Zoo DR1 98,391
MiraBest 1,563
LoTSS DR2 8,805
Figure 3 shows representative, preprocessed examples from each dataset (panels (a)–(c)), illustrating the visual diversity of the three evaluation benchmarks under a shared input pipeline.
Figure 3: Example cutouts from the three evaluation datasets after preprocessing with the same pipeline used for pretraining and evaluation (Section 3). Panels show samples of each class for (a) RGZ DR1, (b) MiraBest, and (c) the LoTSS DR2 visual-classification sample of Horton et al. (2025).
4.1 Common Preprocessing
All evaluation images are treated as single-channel inputs and are standardized using the same per-image ZScale contrast stretch described in Section 3 (mapping to $[0,1]$ and clipping), with NaN/Inf values set to zero. The resulting single-channel image is replicated to three channels and resized to the model input resolution using bicubic interpolation. In downstream evaluation this resolution is model-dependent: ViT-MAE-family models use $224\times 224$ inputs, whereas DINOv2-family models use their native $518\times 518$ inputs. For FITS-based datasets, unreadable files and blank cutouts (e.g., all-zero images) are excluded during indexing.
4.2 MiraBest (Binary FR I/FR II)
MiraBest provides a curated set of morphologically classified radio galaxies intended for machine learning applications (Porter and Scaife, 2023). The released labels include multiple subtypes and confidence levels. For this work we cast MiraBest as a binary Fanaroff–Riley task by mapping all FR I subtypes to FR I and all FR II subtypes to FR II, while excluding hybrid sources. Unless stated otherwise, we further restrict to confident labels and discard uncertain entries to avoid conflating representation quality with label ambiguity. The images are distributed in a batched format; we apply the same normalization and resizing pipeline as for FITS cutouts to obtain a standardized tensor input.
4.3 LoTSS DR2 Visual-Classification Cutouts (Multi-Class Initial Labels)
Horton et al. (2025) present a large visual-classification effort for bright, resolved LoTSS DR2 sources, including an “initial” class taxonomy and additional morphology and environment flags. We download high-resolution LoTSS DR2 cutouts for each source using the public cutout service, centring on the catalog coordinates. Cutout angular size is computed per source from the reported redshift and physical size when available; otherwise we fall back to a fixed cutout size (5 arcmin) and do not apply additional enlargement (factor =1=1).
To keep the benchmark label space well-defined, we use only the initial class taxonomy (FR I, FR II, Hybrid, Spiral, and Relaxed double) and restrict to sources with exactly one active initial label. Samples with no initial label (or ambiguous multiple initial labels) are excluded. We retain the full set of per-source flags in a JSON sidecar for reproducibility and for possible multi-label extensions.
4.4 Radio Galaxy Zoo DR1 (Six-Class Component/Peak Morphology)
Radio Galaxy Zoo DR1 provides citizen-science classifications for complex radio sources, including a consensus level that can be used to filter for high-confidence labels (Wong et al., 2024). From the FIRST subset, we construct a multi-class benchmark based on the discrete component/peak summary provided in the catalog. Specifically, we derive a class label from the number of radio components and peaks and retain the six most common combinations: 1c1p, 1c2p, 1c3p, 2c2p, 2c3p, and 3c3p. We filter the catalog to consensus level $\mathrm{CL}\geq 0.65$ and download FIRST FITS cutouts via the FIRST cutout service, using an adaptive angular size of 1.5 times the catalog largest angular extent (LAE) and enforcing a minimum cutout side length of 8 pixels. To improve data integrity, cutouts are rejected when the WCS-derived image centre deviates from the target coordinate by more than 1 arcmin.
5 ViT Backbone Architecture
This section describes the architectural components used across our continued-pretraining framework. To isolate the effects of radio astronomy-specific data construction, preprocessing, and objective design, we keep the backbone family fixed and explore a small set of controlled modifications: (i) a standard ViT-MAE base encoder, (ii) an MLP projection head used only for the contrastive objective, and (iii) optional register tokens inserted into the encoder token stream.
5.1 Base Encoder: ViT-MAE
STRADAViT builds on the ViT-MAE formulation (He et al., 2022), which uses a Vision Transformer (ViT) encoder (Dosovitskiy et al., 2021; Vaswani et al., 2017) with a lightweight decoder during masked reconstruction. Images are split into non-overlapping $p\times p$ patches and linearly embedded into a sequence of token vectors. A learnable class token is prepended and positional encodings are added, after which the sequence is processed by a stack of Transformer encoder blocks (multi-head self-attention and a feed-forward MLP).
Unless stated otherwise, we use the standard ViT-MAE base configuration (ViT-B/16). We do not modify the encoder embedding dimension, depth, attention head count, or the MAE decoder architecture relative to this default; controlled architectural comparisons are restricted to the optional projection head (contrastive objective only) and optional register tokens (below). We choose this MAE-family starting point because it provides a single, readily modifiable encoder family for both reconstruction and contrastive branches, allowing a controlled study of the transfer-learning training recipe without introducing additional implementation and compute variance from a second pretrained stack. All training views are resized to the model input resolution ($224\times 224$).
In the reconstruction objective, the decoder is used to reconstruct masked patches from the encoder representation. In the contrastive objective, we use the encoder as a feature extractor without the reconstruction decoder. For the feature readout, we use mean pooling over patch tokens (i.e., excluding the class token) to produce a single embedding vector per view. This readout is consistent with the MAE-family objective because the masked-reconstruction objective acts on patch content rather than on a CLS-based representation, so pooled patch tokens provide the more directly optimized transfer feature. Concretely, if $h_{i}\in\mathbb{R}^{D}$ denotes the last-layer embedding of the $i$th patch token (with $i=1,\ldots,N$), we form
$z=\frac{1}{N}\sum_{i=1}^{N}h_{i},$ (1)
and use $z$ as the per-view representation. We do not use the class (CLS) token as a readout in either objective. In the masked reconstruction setting, the objective is defined over masked patch predictions rather than on the CLS token. In the contrastive setting, the loss is applied to $z$ (after the projection head), so the CLS token is not directly optimized as a representation target.
5.2 Projection Head for Contrastive Learning
Contrastive pretraining uses an additional projection head that maps the pooled encoder embedding to a lower-dimensional space in which the contrastive objective is applied. We use a two-layer MLP with hidden dimension 2048 and output dimension 128. The MLP uses LayerNorm after the first linear layer, a ReLU nonlinearity, and Batch Normalization on the output. The resulting vectors are $\ell_{2}$-normalized before computing the contrastive loss. We train on 4 GPUs and synchronize the projection-head BatchNorm statistics across devices (SyncBatchNorm).
5.3 Register Tokens
We also evaluate an optional register-token variant inspired by recent foundation-style ViT designs (Oquab et al., 2023). In this setting, we insert $R$ learnable register tokens immediately after the class token and before the patch tokens. These register tokens participate in all encoder blocks and can act as additional content-bearing slots, while leaving the patch token grid unchanged. For compatibility with masked reconstruction and for consistent downstream readout, register tokens are removed from the sequence before the reconstruction decoder and before patch-token pooling in the contrastive objective. We report results for $R\in\{0,4\}$, where $R=0$ recovers the standard ViT-MAE encoder.
5.4 Continued-Pretraining Branches
We consider three continued-pretraining branches. The reconstruction-only branch applies the masked reconstruction objective by itself. The contrastive-only branch applies the contrastive objective directly from the chosen initial checkpoint, without a preceding reconstruction stage. The two-stage branch first trains with the reconstruction objective and then continues from that checkpoint with the contrastive objective. We use “branch” for these pipelines and reserve “phase 1” and “phase 2” for the two stages of the two-stage branch.
6 Masked Reconstruction Objective
The reconstruction branch trains the encoder–decoder model with a single-view masked reconstruction objective. We first describe the run-time view generation and augmentations used to construct reconstruction inputs, and then detail the reconstruction loss variants used for controlled comparisons against the default ViT-MAE loss.
6.1 View Generation for Masked Reconstruction
The reconstruction branch uses a single-view masked reconstruction objective. Each standardized $512\times 512$ cutout is mapped to one training view with an ROI-aligned strategy intended to reduce empty crops while retaining a controlled fraction of wide-field context. This is important for sparse radio astronomy cutouts, where naive random crops can miss the source and overemphasize background structure.
Given an input cutout, we sample one of two geometric regimes:
  1. 1.
    Wide-field crop (probability $p_{\mathrm{global}}=0.2$): a square random resized crop is taken directly from the full cutout with scale range $[0.70,1.00]$ and resized to the model input size.
  2. 2.
    ROI-aligned crop (probability $1-p_{\mathrm{global}}$): an object-centric square region-of-interest (ROI) is identified by proposing candidate anchors using two complementary heuristics: (i) a $16\times 16$ tile grid is searched for high-significance peaks (top $K_{\mathrm{z}}=3$ candidates, minimum separation of two tiles, and a minimum z-score floor of $2.5$), and (ii) the centres of the top $K_{\mathrm{mean}}=5$ tiles from an $8\times 8$ grid scored by mean absolute intensity are retained as additional candidates. If mean-based candidates are available, we select between the two candidate sets with probability $p_{\mathrm{mean}}=0.5$, and then sample uniformly from the chosen set. Candidate anchors are filtered by requiring their local peak intensity to exceed $0.8$ times the global peak of the parent cutout. A square ROI is then sampled around this anchor with side length fraction in $[0.55,0.80]$ of the cutout size, and a “global-tight” view is sampled inside the ROI with scale range $[0.50,0.70]$ (resized to the model input size), with bounded resampling if an uninformative crop is proposed.
After the geometric step, we apply the mild augmentation regime in Table 3, clamp to $[0,1]$, and then apply the initialization-dependent channel normalization described in Section 8.1. Figure 1 illustrates representative reconstruction-branch views produced by this pipeline.
6.1.1 Augmentation Operators and Parameters
All augmentations operate after ZScale stretching to $[0,1]$. Table 3 summarizes the mild regime used in the reconstruction branch and the strong regime used for the additional contrastive view.
  • Dihedral transforms: random $k\times 90^{\circ}$ rotation with $k\in\{0,1,2,3\}$ and horizontal flip with probability $0.5$.
  • Asinh contrast stretch: random asinh remapping to compress bright structure and lift faint emission.
  • Small affine jitter: bounded translation and isotropic scale perturbation.
  • Additive noise: add pixelwise Gaussian noise with standard deviation $\sigma$.
  • Banding artefacts: random horizontal or vertical bands with additive offset $\Delta$.
  • Flux/background perturbation: multiplicative attenuation of lower-intensity pixels below a threshold derived from the view statistics.
Unless stated otherwise, operators are applied in the order listed above.
Table 3: Augmentation parameter settings used for view generation. “Mild” applies to the reconstruction branch and to the wide view in the contrastive objective; “Strong” applies to the additional contrastive view.
Operator Mild setting Strong setting
Dihedral transforms $k\times 90^{\circ}$ rotation ($k\in\{0,1,2,3\}$) + horizontal flip ($p=0.5$) Same as mild
Asinh stretch $p_{\mathrm{asinh}}=0.35$, $\alpha\in[3,12]$ $p_{\mathrm{asinh}}=0.50$, $\alpha\in[3,20]$
Affine jitter translation $\tau=0.03$, scale jitter $\delta=0.05$ translation $\tau=0.05$, scale jitter $\delta=0.08$
Gaussian noise $\sigma=0.02$ $\sigma=0.04$
Banding reconstruction branch: $B=2$; contrastive wide view: disabled ($B=0$); $\Delta\in[-0.3,0.3]$ $B=3$; $\Delta\in[-0.3,0.3]$
Flux/background perturbation $p=0.5$, $T=\mu+0.3\sigma_{x}$, $f\in[0.7,0.95]$ $p=0.5$, $T=\mu+0.3\sigma_{x}$, $f\in[0.5,0.9]$
6.2 Reconstruction Loss Variants
The reconstruction branch follows the ViT-MAE masked image modeling objective (He et al., 2022). Let $x$ denote an input view, $m$ a binary mask over patch locations, and $\hat{x}$ the model reconstruction (after unpatchifying decoder outputs). The default ViT-MAE training loss is a masked-patch mean squared error (MSE), computed only over the masked patches:
$\mathcal{L}_{\mathrm{MAE}}(x,\hat{x};m)=\frac{1}{|m|}\sum_{j\in m}\lVert\hat{x}_{j}-x_{j}\rVert_{2}^{2}.$ (2)
We treat this as our baseline and compare two explicit modifications that add an additional reconstruction regularizer computed in the same normalized input space as $x$. In the following, $\langle\cdot\rangle$ denotes the mean over all pixels (and channels) of an image.
6.2.1 Baseline: Masked-Patch MSE (ViT-MAE Default)
The baseline objective is
$\mathcal{L}_{\mathrm{L2}}=w_{\mathrm{L2}}\,\mathcal{L}_{\mathrm{MAE}},$ (3)
with weight $w_{\mathrm{L2}}=1.0$.
6.2.2 Masked-Patch MSE + Global L1
To encourage fidelity under an absolute error metric, we add a global (all-pixel) L1 term:
$\mathcal{L}_{\mathrm{L2+L1}}=w_{\mathrm{L2}}\,\mathcal{L}_{\mathrm{MAE}}+w_{\mathrm{L1}}\,\mathcal{L}_{\mathrm{L1}},$ (4)
where $\mathcal{L}_{\mathrm{L1}}=\langle|\hat{x}-x|\rangle$. We use $w_{\mathrm{L2}}=1.0$ and $w_{\mathrm{L1}}=0.1$.
6.2.3 Masked-Patch MSE + Brightness-Weighted L1
As a second variant, we replace the global L1 with a brightness-weighted L1 term that up-weights residuals in brighter regions. Let $w(x)$ be a per-pixel weight proportional to the input brightness and normalized to have unit mean:
$w(x)=\frac{|x|}{\langle|x|\rangle+\epsilon},$ (5)
with $\epsilon>0$ for numerical stability. The objective is
$\mathcal{L}_{\mathrm{L2+BL1}}=w_{\mathrm{L2}}\,\mathcal{L}_{\mathrm{MAE}}+w_{\mathrm{BL1}}\,\mathcal{L}_{\mathrm{BL1}},$ (6)
where $\mathcal{L}_{\mathrm{BL1}}=\langle w(x)\,|\hat{x}-x|\rangle$ and $w(x)$ is normalized so that $\langle w(x)\rangle\approx 1$. We use $w_{\mathrm{L2}}=1.0$ and $w_{\mathrm{BL1}}=0.1$.
7 Contrastive Learning Objective
The contrastive objective is used in two settings: directly in the contrastive-only branch, and as the second stage of the two-stage branch. It learns invariances across multiple augmented views of the same standardized cutout. We apply the contrastive objective to the $\ell_{2}$-normalized projection-head outputs (Section 5), using all other views in the (distributed) batch as negatives. We compare three contrastive loss variants: a standard InfoNCE formulation, a “soft” hard-negative reweighting variant (soft-HCL), and a hard-negative objective (HCL) (Robinson et al., 2021).
7.1 View Generation for Contrastive Learning
The contrastive branch requires correlated views of the same cutout. We therefore use an anchored multi-view strategy that ties all views to a common ROI, reducing false-positive pairs in sparse fields where independent random crops can contain different sources or only background.
For each cutout, we first select an anchor region using the same anchor-selection scheme as in the reconstruction branch and sample an object-centric ROI with side length fraction in $[0.55,0.80]$ of the cutout. In this work we generate two square views (resized to the model input size) with fixed ordering:
  • Wide/global view: sampled within the ROI using scale range $[0.50,0.70]$.
  • Additional view: sampled within the ROI using a narrower scale range $[0.20,0.35]$ and constrained to overlap substantially with the wide view, requiring $\mathrm{IoA}\geq 0.70$, ensuring shared content across views.
We define intersection-over-area (IoA) as
$\mathrm{IoA}(a,b)=\frac{\mathrm{area}(a\cap b)}{\mathrm{area}(a)}.$ (7)
We use IoA rather than a symmetric IoU criterion so that each additional view lies mostly inside the wide view while the wide view can retain broader context. Augmentations follow the same family as in the reconstruction branch, with the strong regime applied to the additional view. Each view is clamped to $[0,1]$ and normalized using the policy in Section 8.1.
7.1.1 Online Non-Empty Resampling
Despite ROI anchoring, some proposed crops remain effectively empty. We therefore use rejection sampling at dataset level: if the reconstruction view or any contrastive view fails a non-empty criterion, the pipeline resamples a different cutout index, up to five attempts. This suppresses uninformative crops without changing the batch interface.
7.2 Contrastive Loss Variants
Each training step operates on a batch of $B$ cutouts, generating $V$ views per cutout (in this work $V=2$) and producing an embedding $z\in\mathbb{R}^{d}$ for each view via the encoder, mean pooling over patch tokens, and the projection head (Section 5). Embeddings are $\ell_{2}$-normalized and concatenated across all devices to form a global set of keys; for each query embedding, positives are the other view(s) of the same cutout and negatives are all remaining keys (except the query itself). Similarities are computed as dot products and scaled by a temperature $\tau$ that is linearly warmed from $0.20$ to $0.10$ over the warmup portion of training (first 15% of epochs) and then kept at $\tau=0.10$. In our distributed implementation, embeddings are all-gathered across devices to form the key set, but embeddings from other devices are treated as stop-gradient keys (i.e., gathered without gradient) while the loss is evaluated for the local queries.
The size of the negative pool is bounded by the distributed micro-batch used to compute the loss (i.e., a single forward pass). Let $G$ denote the number of devices (world size). Each loss evaluation involves $G\!\cdot\!B\!\cdot\!V$ key embeddings (all views across devices). For a given query, the number of negative keys is therefore approximately $G\!\cdot\!B\!\cdot\!V-V$ (excluding the $V$ views of the query’s underlying cutout, including itself). In the runs reported in this work we use $G=4$ GPUs, $B=64$ images per GPU, and $V=2$ views, yielding 256 images (512 keys) per micro-batch and 510 negative keys per query (equivalently, 255 other images). We use gradient accumulation to obtain a fixed effective global batch size of 2048 images per optimizer update. However, gradient accumulation does not increase the negative pool per loss evaluation. InfoNCE, HCL and soft-HCL share the same negative pool; they differ only in how negatives are weighted.
7.2.1 Baseline: InfoNCE
We use a standard temperature-scaled InfoNCE contrastive loss (the NT-Xent form used in SimCLR (Chen et al., 2020)) as our baseline. Let $s_{ij}=z_{i}^{\top}z_{j}$ denote the similarity between a query $z_{i}$ and a key $z_{j}$. Let $P(i)$ be the set of positive keys for query $i$ and $A(i)$ the set of allowed keys in the denominator (all keys excluding the query itself). Let $N(i)=A(i)\setminus P(i)$ denote the set of negatives. Define the positive mass
$p_{i}=\sum_{p\in P(i)}\exp(s_{ip}/\tau),$ (8)
and the per-negative masses $n_{ij}=\exp(s_{ij}/\tau)$ for $j\in N(i)$, with total negative mass
$q_{i}=\sum_{j\in N(i)}n_{ij}.$ (9)
The baseline multi-positive InfoNCE loss can then be written as
$\mathcal{L}_{\mathrm{NCE}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{p_{i}}{p_{i}+q_{i}}.$ (10)
7.2.2 Hard Contrastive Learning (HCL)
We also evaluate hard-negative reweighting following Robinson et al. (2021). HCL keeps the same overall contrastive log-ratio structure as InfoNCE, but it modifies how the negative term is accumulated: negatives that are more similar to the query contribute more to the denominator.
Starting from the baseline denominator $p_{i}+q_{i}$, HCL replaces $q_{i}=\sum_{j\in N(i)}n_{ij}$ with a reweighted negative mass $\tilde{q}^{\mathrm{HCL}}_{i}$ that emphasizes hard negatives. First compute the mean negative mass $\bar{n}_{i}=\mathrm{mean}_{k\in N(i)}(n_{ik})$ and define importance weights
$w^{\mathrm{HCL}}_{ij}=\beta\,\frac{n_{ij}}{\bar{n}_{i}},\qquad j\in N(i),$ (11)
so that “hard” negatives (large $n_{ij}$) receive larger weights. For example, a negative with $n_{ij}=2\bar{n}_{i}$ receives twice the weight of an average negative (for $\beta=1$). The resulting effective negative mass is
$\tilde{q}^{\mathrm{HCL}}_{i}=\sum_{j\in N(i)}w^{\mathrm{HCL}}_{ij}\,n_{ij},$ (12)
which is equivalent to $\tilde{q}^{\mathrm{HCL}}_{i}=\beta\,\frac{\sum_{j\in N(i)}n_{ij}^{2}}{\bar{n}_{i}}$; this makes the denominator disproportionately sensitive to high-similarity negatives through the squared term.
Finally, the HCL loss is
$\mathcal{L}_{\mathrm{HCL}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{p_{i}}{p_{i}+\tilde{q}^{\mathrm{HCL}}_{i}}.$ (13)
The original formulation also permits an optional debiasing correction controlled by $\tau_{+}$; in this work we use $\tau_{+}=0$ (disabled) and set $\beta=1.0$.
7.2.3 Soft Hard-Negative Reweighting (soft-HCL)
Soft-HCL can be viewed as a “softened” alternative to HCL in which negative weights are obtained from a mixture of uniform weights and a hardness distribution computed from (untempered) similarities. Using the same notation as above, we first define a hardness distribution over negatives
$w^{\mathrm{hard}}_{ij}=\frac{\exp(s_{ij}/\tau_{h})}{\sum_{k\in N(i)}\exp(s_{ik}/\tau_{h})},\qquad j\in N(i),$ (14)
and a uniform distribution $w^{\mathrm{uni}}_{ij}=1/|N(i)|$. Note that $\tau_{h}$ is distinct from the contrastive temperature $\tau$; it controls how concentrated the hardness distribution is, independent of the softmax temperature used in the loss. We then form the mixture
$w^{\mathrm{soft}}_{ij}=(1-\alpha)w^{\mathrm{uni}}_{ij}+\alpha w^{\mathrm{hard}}_{ij},$ (15)
scale the weights so that $\sum_{j\in N(i)}w^{\mathrm{soft}}_{ij}=|N(i)|$, and compute an effective negative mass
$\tilde{q}^{\mathrm{soft}}_{i}=\sum_{j\in N(i)}w^{\mathrm{soft}}_{ij}\,n_{ij}.$ (16)
We then use the same loss form as in Eq. (13), replacing $\tilde{q}^{\mathrm{HCL}}_{i}$ with $\tilde{q}^{\mathrm{soft}}_{i}$. We use $\alpha=0.5$ and $\tau_{h}=0.15$.
Intuitively, this modification aims to balance two competing effects: (i) prioritizing informative, “hard” negatives that are most confusable with the query, and (ii) avoiding a setting where a small number of extreme negatives dominate the denominator and destabilize learning. This is relevant in our setting because radio astronomy cutouts are often sparse and background-dominated, so a small subset of negatives can achieve high similarity due to shared noise/artefact structure or repeated morphology, and in rare cases may represent near-duplicates (false negatives) arising from large mosaics and tiling boundaries. The mixing parameter $\alpha$ interpolates between uniform weighting and hardness-based weighting, while $\tau_{h}$ controls how sharply the hard-negative distribution concentrates on the most similar negatives.
Table 4: Contrastive loss variants and hyperparameters used in the contrastive branch.
Variant Parameters
InfoNCE temperature $\tau$: $0.20\rightarrow 0.10$ (warmup 15% epochs), then $\tau=0.10$
HCL $\tau$ as above; $\beta=1.0$, $\tau_{+}=0$
Soft-HCL $\tau$ as above; $\alpha=0.5$, $\tau_{h}=0.15$
8 Training Setup
This section summarizes (i) the initialization and input-normalization choices and (ii) the optimization settings and training schedules used in the reconstruction and contrastive branches.
8.1 Initialization and Input Normalization
We consider two starting points for continued pretraining: (i) continued pretraining from publicly available ViT-MAE weights pretrained on ImageNet-1K, and (ii) scratch training with random initialization. The choice of starting point also determines whether we apply an additional channel-wise input normalization step. When continuing from ImageNet-pretrained weights, we apply the standard ImageNet mean and standard deviation normalization to match the pretrained backbone. When training from scratch, we omit this normalization (identity), since our preprocessing already maps inputs to $[0,1]$. The contrastive-only branch initializes directly from the chosen starting checkpoint, whereas the two-stage branch initializes its contrastive stage from the corresponding reconstruction checkpoint trained with the same architectural settings.
8.2 Optimization and Training Hyperparameters
Unless stated otherwise, settings are shared across runs and the effective global batch size is fixed at 2048 images per optimizer update via gradient accumulation.
8.2.1 Common Optimization Settings
[…]
8.2.2 Distributed Training and Normalization Details
Training is performed using data-parallel distributed training across GPUs. In the contrastive phase, each GPU produces a local set of normalized embeddings for all views in its micro-batch; we then all-gather these embeddings across devices to form a global key set for the contrastive denominator. The loss is evaluated for the local queries using this global key set, with remote embeddings treated as stop-gradient keys (gathered without gradient). This design keeps the negative pool consistent across devices while avoiding the memory overhead of backpropagating through cross-device keys. Because the projection head includes a BatchNorm layer, we use SyncBatchNorm across the 4 GPUs to ensure consistent normalization statistics.
8.2.3 Reconstruction-Branch Training
Unless stated otherwise, the reconstruction branch trains for 35 epochs with a fixed mask ratio $0.75$ and patch size $16$. We use per-device batch size 256 and apply gradient clipping with $\lVert g\rVert_{2}$ capped at 1.0. We compare the three reconstruction losses described in Section 6 (masked-patch MSE; MSE + L1; MSE + brightness-weighted L1), using the weights specified there.
8.2.4 Contrastive-Branch Training
Unless stated otherwise, the contrastive branch trains for 35 epochs, both when used directly in the contrastive-only branch and as the second stage of the two-stage branch. We generate $V=2$ views per cutout and use per-device batch size 64 (so each step processes $2B$ views). The encoder is unmasked in this branch (no patch masking), and we apply gradient clipping with $\lVert g\rVert_{2}$ capped at 3.0. The contrastive temperature is warmed linearly from $\tau=0.20$ to $\tau=0.10$ over the warmup portion of training (first 15% of epochs) and then kept at $\tau=0.10$.
[…]
The longer token sequence also reduces batch size. ViT-MAE-initialized contrastive runs use per-device batch size 64, whereas DINOv2-initialized runs use 20. Gradient accumulation keeps the effective image batch near 2048 in both cases, but the contrastive denominator is set by the distributed micro-batch. With 4 GPUs and $V=2$ views, the per-step key set therefore drops from 512 views to 160 views, reducing negatives per query from approximately 510 to 158. The DINOv2-initialized runs are therefore interpreted as a targeted initialization study rather than as a perfectly matched repeat of the ViT-MAE experiments.
Table 5 provides a compact summary of the default branch-specific settings.
Table 5: Summary of the default training hyperparameters used in the reconstruction and contrastive branches.
Setting Reconstruction branch Contrastive branch
Epochs 35 35
Patch size 16 16
Masking mask ratio $0.75$ unmasked encoder (no patch masking)
Views per cutout 1 $V=2$
Per-device batch size 256 64
Effective global batch 2048 images via gradient accumulation (before view replication)
Optimizer AdamW (fused on CUDA), weight decay $0.05$, $(\beta_{1},\beta_{2})=(0.9,0.95)$
Learning-rate schedule cosine decay with linear warmup (15% of epochs), base LR $5\times 10^{-4}$
Precision bf16
Gradient clipping $\lVert g\rVert_{2}\leq 1.0$ $\lVert g\rVert_{2}\leq 3.0$
Temperature n/a $\tau:0.20\rightarrow 0.10$ (warmup), then $\tau=0.10$
9 Transfer Evaluation Protocol
This section describes our downstream transfer evaluation protocol, covering both linear probing and full fine-tuning. Both evaluations follow the same data preprocessing, cross-validation, and metric reporting, and differ only in which model parameters are optimized.
9.1 Common Evaluation Setup
  1. 1.
    Preprocessing and resizing: All evaluation images are standardized as described in Section 3 (per-image ZScale contrast stretch to $[0,1]$), replicated to three channels, and resized to the model-specific input resolution. For ViT-MAE-family checkpoints this is $224\times 224$, whereas for DINOv2-family checkpoints we retain the native $518\times 518$ resolution specified by the checkpoint configuration or image processor.
  2. 2.
    Augmentations: During training (both linear probe and fine-tuning), we apply only simple dihedral (D4) transforms: a random rotation by $k\times 90^{\circ}$ with $k\in\{0,1,2,3\}$ and a horizontal flip with probability 0.5 (implemented with torch.rot90 and torch.flip, so no interpolation is introduced). Evaluation uses no stochastic augmentations.
  3. 3.
    Input normalization: Channel-wise normalization uses the mean and standard deviation stored with each checkpoint’s image processor: ImageNet mean/std for ImageNet-initialized checkpoints and identity normalization for scratch-trained checkpoints (Section 8.1).
  4. 4.
    Cross-validation: For each benchmark dataset independently, we run $K=3$-fold stratified cross-validation using a fixed random seed (0) with shuffling enabled. In each fold, models are trained on K1K-1 splits, which amounts to approximately $2/3$ of the data for training (about 66.7%), and evaluated on the held-out split, which amounts to approximately $1/3$ of the data for testing/evaluation (about 33.3%). We do not use an additional validation split or early stopping; models are trained for a fixed number of epochs and evaluated on the held-out split once per epoch, with final reported metrics computed on the held-out split after training. Reported results are the mean and standard deviation across folds. To mitigate class imbalance, we use a class-weighted cross-entropy loss with “balanced” weights computed from the dataset class frequencies.
  5. 5.
    Metrics: We report macro-averaged and weighted-averaged F1 scores. Macro-F1 assigns equal weight to each class, while weighted-F1 weights per-class F1 by class support. We also aggregate a confusion matrix by summing fold confusion matrices. These confusion-matrix diagnostics are used in Section 10.3 to compare the selected STRADAViT model with the two starting-point baselines.
9.2 Linear Probing
We evaluate representation quality with a strict linear probe. For an input image, we obtain the encoder last-layer token embeddings and form a single feature vector by mean pooling over patch tokens (excluding the CLS token), consistent with the contrastive-branch readout used for MAE-family models:
$z=\frac{1}{N}\sum_{i=1}^{N}h_{i}.$ (17)
We then train a linear classifier $W\in\mathbb{R}^{C\times D}$ with logits $\ell=Wz+b$, where $C$ is the number of classes and $D$ is the encoder hidden dimension. “Strict” denotes that (i) all encoder parameters are frozen ($\nabla_{\theta_{\mathrm{enc}}}=0$), (ii) the probe head is linear-only (any additional normalization, dropout, or MLP components in the classification head are removed/disabled), and (iii) the frozen backbone is kept in evaluation mode during probe training so that stochastic layers do not introduce additional variation. Backbone-internal normalizations (e.g., any frozen pooling/normalization that is part of the backbone) remain part of the representation, but no additional trainable layers are introduced beyond the linear classifier.
We train the probe for 20 epochs using AdamW with weight decay $0.05$ and a cosine learning-rate schedule with warmup ratio 0.05. Since only the linear classifier is trainable, we use a single learning rate for the probe head ($5\times 10^{-3}$) and clip gradients to $\lVert g\rVert_{2}\leq 1.0$. Training uses bf16 mixed precision and is run with data-parallel training on 4 GPUs (Section 8); for linear probing we use an effective global batch size of 256 images.
9.3 Fine-Tuning
In fine-tuning, we train the full model end-to-end: all encoder parameters and the classification head are optimized on the downstream label space. For MAE-family models, our classification head is lightweight and batch-size agnostic: LayerNorm, optional dropout, and a single linear layer (no BatchNorm). We fine-tune for 20 epochs using AdamW with weight decay $0.05$ and a cosine learning-rate schedule with warmup ratio 0.2. We use a base learning rate of $5\times 10^{-4}$ for the backbone and a head learning-rate multiplier of 10.0. Following MAE-style transfer practice, we apply layer-wise learning-rate decay with factor 0.65 across the Transformer blocks, such that earlier layers receive smaller learning rates than later layers. We clip gradients to $\lVert g\rVert_{2}\leq 1.0$ and train with bf16 mixed precision on 4 GPUs (Section 8), using an effective global batch size of 256 images.
10 Results and Analysis
We report transfer results under the evaluation protocol described in Section 9. Results are organized into baseline comparisons, branch-wise continued-pretraining results (reconstruction-only, contrastive-only, and two-stage), a targeted DINOv2 initialization ablation under fixed HCL training, a classwise comparison of the selected STRADAViT configuration against the two starting-point baselines, and an overall interpretation of the branch comparison.
Throughout, we report both linear probing and full fine-tuning because they probe different aspects of transfer. Linear probing remains a standard SSL diagnostic for frozen representation quality, and it is especially relevant when encoders are reused as generic backbones or when labeled data are limited. In our setting, improvements under a frozen-backbone probe can be larger and more consistent than improvements after end-to-end fine-tuning, where supervised adaptation can partially close gaps between pretraining variants. We therefore interpret probe gains as evidence of improved linear separability of the learned representations—a desirable property for transfer-oriented reuse—while treating fine-tuning as the primary measure of practical downstream performance on these benchmarks. All baselines and continued-pretraining variants are evaluated with the same downstream preprocessing, folds, and optimization protocol so that these comparisons remain controlled.
10.1 Baselines (DINOv2, ViT-MAE)
We evaluate publicly available baseline checkpoints without any radio astronomy pretraining, using the same linear-probe and fine-tuning procedures (Section 9). In all results tables, model names with the suffix “(R)” denote variants that use register tokens (Section 5). For DINOv2 baselines we use the backbone’s native pooled representation (global pooling with the checkpoint’s built-in normalization, when enabled) and native $518\times 518$ inputs, while for MAE-family models we use patch-token mean pooling with the CLS token excluded and $224\times 224$ inputs.
Tables 6 and 7 summarize baseline performance across datasets using Macro-F1 and Weighted-F1 (mean $\pm$ standard deviation over $K=3$ folds).
Table 6: Baseline model performance under linear probing (mean $\pm$ std over $K=3$ folds). “(R)” indicates register tokens. Best Macro-F1 per dataset is highlighted in bold; ties are broken by Weighted-F1.
Model MiraBest LoTSS DR2 RGZ DR1
Macro-F1 Weighted-F1 Macro-F1 Weighted-F1 Macro-F1 Weighted-F1
DINOv2-Base $0.717\pm 0.014$ $0.718\pm 0.014$ $0.569\pm 0.005$ $0.751\pm 0.004$ $0.661\pm 0.004$ $0.871\pm 0.001$
DINOv2-Small $0.713\pm 0.018$ $0.713\pm 0.018$ $0.544\pm 0.014$ $0.729\pm 0.007$ $0.648\pm 0.002$ $0.861\pm 0.001$
DINOv2-Base(R) $0.685\pm 0.013$ $0.686\pm 0.013$ $0.529\pm 0.009$ $0.716\pm 0.003$ $0.629\pm 0.001$ $0.862\pm 0.001$
ViT-MAE-Base $0.644\pm 0.020$ $0.645\pm 0.020$ $0.489\pm 0.005$ $0.663\pm 0.006$ $0.586\pm 0.001$ $0.841\pm 0.000$
Table 7: Baseline performance under full fine-tuning (mean $\pm$ std over $K=3$ folds). “(R)” indicates register tokens. Best Macro-F1 per dataset is highlighted in bold; ties are broken by Weighted-F1.
Model MiraBest LoTSS DR2 RGZ DR1
Macro-F1 Weighted-F1 Macro-F1 Weighted-F1 Macro-F1 Weighted-F1
DINOv2-Base $0.461\pm 0.195$ $0.451\pm 0.202$ $0.708\pm 0.008$ $31\pm 0.002 0.808$$08\pm 0.003$ $0.916\pm 0.001$
DINOv2-Small $0.707\pm 0.031$ $0.708\pm 0.031$ $0.704\pm 0.014$ $0.828\pm 0.004$ $0.795\pm 0.002$ $0.912\pm 0.001$
DINOv2-Base(R) $0.704\pm 0.102$ $0.704\pm 0.102$ $0.708\pm 0.017$ $0.829\pm 0.004$ $0.812\pm 0.002$ $0.917\pm 0.001$
ViT-MAE-Base $0.724\pm 0.002$ $0.725\pm 0.002$ $0.696\pm 0.015$ $0.819\pm 0.001$ $0.805\pm 0.003$ $0.916\pm 0.001$
Under linear probing (Table 6), DINOv2 remains stronger than ViT-MAE. DINOv2-Base is the best baseline on all three datasets: Macro-F1 reaches $0.717\pm 0.014$ on MiraBest, $0.569\pm 0.005$ on LoTSS DR2, and $0.661\pm 0.004$ on RGZ DR1. DINOv2-Small remains consistently close behind. By contrast, the register-token DINOv2 variant is weaker on every benchmark in the frozen-feature setting, with the largest drop on RGZ DR1 ($0.661\pm 0.004\rightarrow 0.629\pm 0.001$). The strongest off-the-shelf starting point in terms of raw linear separability is therefore the non-register DINOv2 backbone.
Fine-tuning yields a less uniform ranking (Table 7). The strongest baseline is ViT-MAE-Base on MiraBest ($0.724\pm 0.002$), DINOv2-Base on LoTSS DR2 ($0.708\pm 0.008$), and DINOv2-Base(R) on RGZ DR1 ($0.812\pm 0.002$). On LoTSS DR2, DINOv2-Base is retained over DINOv2-Base(R) after a Macro-F1 tie at $0.708$ because it has the higher Weighted-F1. The register variant therefore no longer dominates the fine-tuning baselines, although it remains the strongest DINOv2 model on RGZ DR1. Optimization sensitivity also remains pronounced: the non-register DINOv2-Base variant is unstable on MiraBest ($0.461\pm 0.195$), whereas DINOv2-Small ($0.707\pm 0.031$), DINOv2-Base(R) ($0.704\pm 0.102$), and ViT-MAE-Base ($0.724\pm 0.002$) are more stable there. The fine-tuning ranking is therefore more sensitive to dataset and configuration than the linear-probe ranking.
Across LoTSS DR2 and RGZ DR1, Weighted-F1 remains substantially higher than Macro-F1. Stronger overall performance can therefore coexist with weaker minority-class behavior. These baselines provide the reference point for the continued-pretraining comparisons below.
10.2 Continued Pretraining Branches
10.2.1 Reconstruction-Only Branch
In this section we evaluate continued pretraining runs initialized from ImageNet-pretrained ViT-MAE weights (Section 8.1). These phase 1 results compare the reconstruction losses from Section 6 while keeping the rest of the protocol fixed.
Tables 8 and 9 summarize downstream transfer performance for the reconstruction branch, stratified by the presence of register tokens ($R\in\{0,4\}$).
Table 8: Continued pretraining (reconstruction-only branch): linear-probe results (mean $\pm$ std over $K=3$ folds). Best Macro-F1 per dataset within each register setting is highlighted in bold; ties are broken by Weighted-F1.
Reconstruction loss MiraBest LoTSS DR2 RGZ DR1
Macro-F1 Weighted-F1 Macro-F1 Weighted-F1 Macro-F1 Weighted-F1
No registers ($R=0$)
L2 $0.607\pm 0.013$ $0.607\pm 0.013$ $0.433\pm 0.006$ $0.606\pm 0.005$ $0.526\pm 0.005$ $0.811\pm 0.002$
L2+L1 $0.609\pm 0.006$ $0.609\pm 0.006$ $0.436\pm 0.013$ $0.599\pm 0.009$ $0.525\pm 0.001$ $0.813\pm 0.001$
L2+BL1 $0.609\pm 0.012$ $0.610\pm 0.012$ $0.448\pm 0.008$ $0.611\pm 0.008$ $0.526\pm 0.002$ $0.813\pm 0.001$
With registers ($R=4$)
L2 $0.612\pm 0.012$ $0.613\pm 0.012$ $0.485\pm 0.017$ $0.659\pm 0.014$ $0.565\pm 0.001$ $0.832\pm 0.001$
L2+L1 $5\pm 0.009 0.606±0.0$$6\pm 0.009$ $0.479\pm 0.009$ $0.658\pm 0.013$ $0.562\pm 0.004$ $0.831\pm 0.002$
L2+BL1 $0.608\pm 0.013$ $0.610\pm 0.013$ $0.478\pm 0.013$ $0.660\pm 0.010$ $0.560\pm 0.003$ $0.830\pm 0.001$
Table 8 shows a modest but consistent register effect under linear probing, especially on LoTSS DR2 and RGZ DR1. By contrast, the loss choice has limited impact: without registers, L2+BL1 is selected after tie-breaking, whereas with registers plain L2 is strongest across datasets.
Table 9: Continued pretraining (reconstruction-only branch): full fine-tuning results (mean $\pm$ std over $K=3$ folds). Best Macro-F1 per dataset within each register setting is highlighted in bold; ties are broken by Weighted-F1.
Reconstruction loss MiraBest LoTSS DR2 RGZ DR1
Macro-F1 Weighted-F1 Macro-F1 Weighted-F1 Macro-F1 Weighted-F1
No registers ($R=0$)
L2 $0.708\pm 0.011$ $0.709\pm 0.011$ $0.661\pm 0.013$ $0.791\pm 0.003$ $0.795\pm 0.005$ $0.915\pm 0.001$
L2+L1 $0.713\pm 0.018$ $0.713\pm 0.018$ $0.657\pm 0.017$ $0.793\pm 0.003$ $0.795\pm 0.005$ $0.914\pm 0.001$
L2+BL1 $0.708\pm 0.025$ $0.709\pm 0.026$ $0.654\pm 0.013$ $0.788\pm 0.003$ $0.795\pm 0.005$ $0.914\pm 0.001$
With registers ($R=4$)
L2 $0.717\pm 0.014$ $0.718\pm 0.014$ $0.664\pm 0.015$ $0.792\pm 0.006$ $0.793\pm 0.004$ $0.914\pm 0.001$
L2+L1 $0.703\pm 0.018$ $0.703\pm 0.018$ $0.647\pm 0.021$ $0.785\pm 0.010$ $0.793\pm 0.003$ $0.913\pm 0.000$
L2+BL1 $0.697\pm 0.007$ $0.697\pm 0.007$ $0.655\pm 0.026$ $0.787\pm 0.008$ $0.791\pm 0.006$ $0.913\pm 0.002$
Under full fine-tuning (Table 9), the register advantage largely disappears and the loss choice again has limited impact. The default L2 objective remains competitive or best in most settings, and the entire reconstruction branch occupies a narrow performance band with comparatively small standard deviations. Reconstruction alone is therefore a weak transfer mechanism in this study, even though it behaves more predictably than the most volatile DINO baselines.
10.2.2 Contrastive-Only Branch
We also evaluate the contrastive-only branch, in which the contrastive objective is applied directly from the chosen initialization checkpoint without a preceding reconstruction stage. We compare InfoNCE, Soft-HCL, and HCL under the same downstream protocol for the ViT-MAE-based runs. The separate DINOv2 initialization study is reported below under a fixed HCL recipe.
Tables 10 and 11 summarize the contrastive-only results for the ViT-MAE-based runs, stratified by the presence of register tokens ($R\in\{0,4\}$).
Table 10: Continued pretraining (contrastive-only branch): linear-probe results for the ViT-MAE-based runs (mean $\pm$ std over $K=3$ folds). For each register setting, best Macro-F1 per dataset is highlighted in bold; ties are broken by Weighted-F1.
Contrastive loss MiraBest LoTSS DR2 RGZ DR1
Macro-F1 Weighted-F1 Macro-F1 Weighted-F1 Macro-F1 Weighted-F1
ViT-MAE init, no registers ($R=0$)
InfoNCE $0.676\pm 0.003$ $0.677\pm 0.002$ $0.538\pm 0.010$ $0.710\pm 0.007$ $0.699\pm 0.001$ $0.878\pm 0.000$
Soft-HCL $0.692\pm 0.012$ $0.693\pm 0.012$ $0.549\pm 0.009$ $0.719\pm 0.003$ $0.709\pm 0.001$ $0.883\pm 0.001$
HCL $0.686\pm 0.010$ $0.687\pm 0.011$ $0.548\pm 0.008$ $0.718\pm 0.002$ $0.713\pm 0.004$ $0.886\pm 0.000$
ViT-MAE init, with registers ($R=4$)
InfoNCE $0.662\pm 0.028$ $0.663\pm 0.029$ $0.558\pm 0.023$ $0.730\pm 0.010$ $0.724\pm 0.002$ $0.893\pm 0.000$
Soft-HCL $0.691\pm 0.010$ $0.692\pm 0.009$ $0.560\pm 0.012$ $0.727\pm 0.008$ $0.728\pm 0.003$ $0.896\pm 0.000$
HCL $0.678\pm 0.014$ $0.679\pm 0.015$ $0.560\pm 0.007$ $0.728\pm 0.003$ $0.727\pm 0.003$ $0.895\pm 0.001$
Table 11: Continued pretraining (contrastive-only branch): full fine-tuning results for the ViT-MAE-based runs (mean $\pm$ std over $K=3$ folds). For each register setting, best Macro-F1 per dataset is highlighted in bold; ties are broken by Weighted-F1.
Contrastive loss MiraBest LoTSS DR2 RGZ DR1
Macro-F1 Weighted-F1 Macro-F1 Weighted-F1 Macro-F1 Weighted-F1
ViT-MAE init, no registers ($R=0$)
InfoNCE $0.726\pm 0.005$ $0.727\pm 0.005$ $0.672\pm 0.011$ $0.798\pm 0.004$ $0.818\pm 0.003$ $0.921\pm 0.001$
Soft-HCL $0.719\pm 0.005$ $0.720\pm 0.015$ $0.660\pm 0.002$ $0.799\pm 0.003$ $0.824\pm 0.003$ $0.923\pm 0.001$
HCL $0.723\pm 0.008$ $0.724\pm 0.007$ $0.661\pm 0.005$ $0.798\pm 0.004$ $0.825\pm 0.003$ $0.924\pm 0.001$
ViT-MAE init, with registers ($R=4$)
InfoNCE $0.726\pm 0.010$ $0.727\pm 0.010$ $0.666\pm 0.011$ $0.795\pm 0.004$ $0.819\pm 0.002$ $0.922\pm 0.001$
Soft-HCL $0.735\pm 0.014$ $0.736\pm 0.013$ $0.667\pm 0.007$ $0.798\pm 0.003$ $0.822\pm 0.003$ $0.923\pm 0.002$
HCL $0.722\pm 0.013$ $0.723\pm 0.013$ $0.665\pm 0.003$ $0.797\pm 0.003$ $0.826\pm 0.001$ $0.924\pm 0.001$
The contrastive branch improves substantially on the phase 1 branch under both linear probing and fine-tuning. The main gain therefore comes from the contrastive objective rather than from masked reconstruction. Loss rankings remain dataset-dependent: Soft-HCL is strongest in several MiraBest and LoTSS DR2 settings, whereas HCL is the most reliable choice on RGZ DR1. Registers interact with the loss in the same dataset-dependent way and do not produce a uniform ordering across datasets.
10.2.3 DINOv2 Initialization Ablation
To test portability of the same adaptation pipeline to a differently pretrained ViT family, we also perform a targeted DINOv2 initialization ablation based on fixed HCL training. This provides a direct check of whether the radio astronomy adaptation pipeline transfers to a backbone whose original pretraining is already contrastive and more structured than ViT-MAE. Table 12 summarizes this ablation.
Table 12: DINOv2 initialization ablation under the contrastive-only HCL recipe (mean $\pm$ std over $K=3$ folds).
Setting Initialization MiraBest LoTSS DR2 RGZ DR1
Macro-F1 Weighted-F1 Macro-F1 Weighted-F1 Macro-F1 Weighted-F1
Linear Probe DINOv2-Base
Linear Probe DINOv2-Base(R) $0.686\pm 0.018$ $0.687\pm 0.018$ $0.561\pm 0.005$ $0.731\pm 0.003$ $0.739\pm 0.004$ $0.895\pm 0.000$
Fine-Tuning DINOv2-Base
Fine-Tuning DINOv2-Base(R) $0.738\pm 0.014$ $0.738\pm 0.014$ $0.678\pm 0.015$ $0.816\pm 0.006$ $0.828\pm 0.002$ $0.922\pm 0.002$
The DINOv2-based results reported here show that the same contrastive adaptation recipe is not specific to the ViT-MAE starting point. Under linear probing, the DINOv2-Base(R)-initialized run improves on the corresponding off-the-shelf DINOv2-Base(R) baseline on all three datasets, with the largest change on RGZ DR1. Under full fine-tuning, the adapted model improves on the off-the-shelf DINOv2-Base(R) baseline on MiraBest and RGZ DR1, but not on LoTSS DR2. Relative to the ViT-MAE-based adapted models, the DINOv2-initialized run is broadly competitive, with only small differences across all three datasets. The main exception remains LoTSS DR2, where the adapted DINOv2 result follows the same pattern as the ViT-MAE-based adapted models and remains below the off-the-shelf baselines.
10.2.4 Two-Stage Branch
The two-stage branch initializes its second stage from the corresponding reconstruction checkpoint and then refines the encoder with a contrastive objective (Section 7). We compare contrastive loss variants while holding the reconstruction-stage recipe fixed.
Tables 13 and 14 report two-stage transfer results for both register settings.
Table 13: Continued pretraining (two-stage branch): linear-probe results (mean $\pm$ std over $K=3$ folds). The contrastive stage uses an unmasked encoder. Best Macro-F1 per dataset within each register setting is highlighted in bold; ties are broken by Weighted-F1.
Reconstruction loss Contrastive loss MiraBest LoTSS DR2 RGZ DR1
Macro-F1 Weighted-F1 Macro-F1 Weighted-F1 Macro-F1 Weighted-F1
No registers ($R=0$)
L2 InfoNCE $0.670\pm 0.010$ $0.671\pm 0.010$ $0.529\pm 0.010$ $0.705\pm 0.006$ $0.693\pm 0.003$ $0.877\pm 0.001$
Soft-HCL $0.677\pm 0.007$ $0.678\pm 0.007$ $0.543\pm 0.010$ $0.715\pm 0.004$ $0.704\pm 0.002$ $0.881\pm 0.000$
HCL $0.689\pm 0.010$ $0.690\pm 0.011$ $0.545\pm 0.012$ $0.714\pm 0.006$ $0.706\pm 0.003$ $0.884\pm 0.001$
L2+L1 InfoNCE $0.664\pm 0.013$ $0.665\pm 0.013$ $0.528\pm 0.010$ $0.701\pm 0.006$ $0.693\pm 0.004$ $0.876\pm 0.001$
Soft-HCL $0.678\pm 0.004$ $0.679\pm 0.004$ $0.564\pm 0.007$ $0.739\pm 0.005$ $0.701\pm 0.002$ $0.881\pm 0.000$
HCL $0.678\pm 0.008$ $0.679\pm 0.007$ $0.547\pm 0.015$ $17\pm 0.007 0.707$$07\pm 0.005$ $0.884\pm 0.001$
L2+BL1 InfoNCE $0.671\pm 0.015$ $0.673\pm 0.016$ $0.535\pm 0.012$ $0.706\pm 0.006$ $0.690\pm 0.003$ $0.876\pm 0.000$
Soft-HCL $0.680\pm 0.001$ $0.681\pm 0.001$ $0.550\pm 0.014$ $0.722\pm 0.008$ $0.706\pm 0.002$ $0.883\pm 0.000$
HCL $0.697\pm 0.013$ $0.698\pm 0.014$ $0.552\pm 0.011$ $0.717\pm 0.004$ $0.708\pm 0.003$ $0.884\pm 0.001$
With registers ($R=4$)
L2 InfoNCE $0.671\pm 0.027$ $0.671\pm 0.027$ $0.566\pm 0.009$ $0.735\pm 0.004$ $0.722\pm 0.002$ $0.892\pm 0.001$
Soft-HCL $0.672\pm 0.011$ $0.673\pm 0.012$ $0.575\pm 0.005$ $0.738\pm 0.005$ $0.727\pm 0.003$ $0.893\pm 0.000$
HCL $0.693\pm 0.001$ $0.694\pm 0.001$ $0.571\pm 0.009$ $0.743\pm 0.006$ $0.731\pm 0.004$ $0.894\pm 0.001$
L2+L1 InfoNCE $0.675\pm 0.017$ $0.675\pm 0.018$ $0.566\pm 0.014$ $0.731\pm 0.005$ $0.718\pm 0.002$ $0.890\pm 0.001$
Soft-HCL $0.674\pm 0.023$ $0.675\pm 0.023$ $0.564\pm 0.007$ $0.739\pm 0.005$ $0.732\pm 0.002$ $0.894\pm 0.002$
HCL $0.699\pm 0.016$ $0.700\pm 0.016$ $0.565\pm 0.012$ $0.738\pm 0.007$ $0.730\pm 0.004$ $0.894\pm 0.001$
L2+BL1 InfoNCE $0.672\pm 0.019$ $0.673\pm 0.020$ $0.559\pm 0.012$ $0.730\pm 0.004$ $0.719\pm 0.003$ $0.891\pm 0.001$
Soft-HCL $0.680\pm 0.010$ $0.680\pm 0.010$ $0.558\pm 0.009$ $0.730\pm 0.003$ $0.729\pm 0.002$ $0.894\pm 0.001$
HCL $0.668\pm 0.004$ $0.669\pm 0.005$ $0.565\pm 0.013$ $0.737\pm 0.007$ $0.730\pm 0.001$ $0.894\pm 0.001$
Table 14: Continued pretraining (two-stage branch): full fine-tuning results (mean $\pm$ std over $K=3$ folds). The contrastive stage uses an unmasked encoder. Best Macro-F1 per dataset within each register setting is highlighted in bold; ties are broken by Weighted-F1.
Reconstruction loss Contrastive loss MiraBest LoTSS DR2 RGZ DR1
Macro-F1 Weighted-F1 Macro-F1 Weighted-F1 Macro-F1 Weighted-F1
No registers ($R=0$)
L2 InfoNCE $0.714\pm 0.002$ $0.715\pm 0.015$ $0.664\pm 0.013$ $0.793\pm 0.004$ $0.824\pm 0.003$ $0.922\pm 0.001$
Soft-HCL $0.732\pm 0.007$ $0.733\pm 0.007$ $0.661\pm 0.007$ $0.790\pm 0.004$ $0.822\pm 0.003$ $0.923\pm 0.001$
HCL $0.727\pm 0.003$ $0.728\pm 0.003$ $0.654\pm 0.008$ $0.793\pm 0.004$ $0.824\pm 0.004$ $0.923\pm 0.002$
L2+L1 InfoNCE $0.715\pm 0.009$ $0.716\pm 0.009$ $0.655\pm 0.006$ $0.790\pm 0.003$ $0.821\pm 0.003$ $0.922\pm 0.001$
Soft-HCL $0.723\pm 0.018$ $0.724\pm 0.018$ $0.654\pm 0.003$ $0.791\pm 0.004$ $0.821\pm 0.002$ $0.923\pm 0.001$
HCL $0.719\pm 0.001$ $0.720\pm 0.000$ $0.656\pm 0.006$ $0.791\pm 0.005$ $0.825\pm 0.001$ $0.924\pm 0.001$
L2+BL1 InfoNCE $0.725\pm 0.014$ $0.726\pm 0.015$ $0.649\pm 0.004$ $0.790\pm 0.005$ $0.821\pm 0.003$ $0.922\pm 0.001$
Soft-HCL $0.725\pm 0.011$ $0.726\pm 0.011$ $0.664\pm 0.004$ $0.797\pm 0.002$ $0.823\pm 0.002$ $0.922\pm 0.001$
HCL $0.733\pm 0.011$ $0.733\pm 0.010$ $0.641\pm 0.011$ $0.794\pm 0.002$ $0.821\pm 0.004$ $0.923\pm 0.002$
With registers ($R=4$)
L2 InfoNCE $0.726\pm 0.006$ $0.727\pm 0.006$ $0.667\pm 0.013$ $0.795\pm 0.007$ $0.817\pm 0.002$ $0.921\pm 0.001$
Soft-HCL $0.736\pm 0.014$ $0.737\pm 0.014$ $0.665\pm 0.005$ $0.799\pm 0.004$ $0.824\pm 0.002$ $0.924\pm 0.001$
HCL $0.730\pm 0.009$ $0.731\pm 0.009$ $0.665\pm 0.004$ $0.803\pm 0.001$ $0.824\pm 0.004$ $0.924\pm 0.001$
L2+L1 InfoNCE $0.714\pm 0.015$ $0.715\pm 0.015$ $0.649\pm 0.006$ $0.792\pm 0.004$ $0.820\pm 0.001$ $0.921\pm 0.001$
Soft-HCL $0.725\pm 0.004$ $0.726\pm 0.011$ $0.677\pm 0.014$ $0.803\pm 0.008$ $0.825\pm 0.001$ $0.924\pm 0.001$
HCL $0.721\pm 0.007$ $0.722\pm 0.007$ $0.664\pm 0.006$ $0.801\pm 0.003$ $0.822\pm 0.006$ $0.923\pm 0.001$
L2+BL1 InfoNCE $0.734\pm 0.010$ $0.734\pm 0.010$ $0.659\pm 0.015$ $0.793\pm 0.007$ $0.819\pm 0.002$ $0.921\pm 0.001$
Soft-HCL $0.732\pm 0.008$ $0.732\pm 0.008$ $0.654\pm 0.010$ $0.795\pm 0.005$ $0.822\pm 0.001$ $0.922\pm 0.001$
HCL $0.727\pm 0.009$ $0.728\pm 0.009$ $0.655\pm 0.006$ $0.798\pm 0.002$ $0.822\pm 0.001$ $0.923\pm 0.001$
Including the contrastive-only results clarifies the role of the reconstruction stage. Under linear probing, the best two-stage models remain the strongest STRADAViT configurations overall, but the extra reconstruction stage helps mainly on MiraBest and LoTSS DR2; RGZ DR1 is already well served by contrastive-only training. Relative to the strongest off-the-shelf DINOv2 baseline, the best two-stage probe result remains below DINOv2 on MiraBest but exceeds it on LoTSS DR2 and RGZ DR1.
The loss ranking inside the two-stage branch remains dataset-dependent. InfoNCE is useful as a reference, but the strongest probe settings almost always involve HCL or Soft-HCL. Under full fine-tuning, phase 1 again helps most on MiraBest and LoTSS DR2, while RGZ DR1 remains effectively saturated by contrastive-only training. Even after two-stage continued pretraining, strong off-the-shelf baselines remain demanding comparators, especially on LoTSS DR2.
10.3 Selected STRADAViT Model: Summary and Classwise Analysis
Table 15: Macro-F1 percentage-point ($p$) changes for the best reported two-stage model on each dataset, relative to the ViT-MAE starting point used for continued pretraining and to the strongest DINOv2 variant. Linear-probe and fine-tuning settings are reported separately. Positive values indicate improvement.
Setting Reference MiraBest LoTSS DR2 RGZ DR1
Linear Probe
vs ViT-MAE +5.5+5.5 +8.6+8.6 +14.6+14.6
vs strongest DINOv2 1.8-1.8 +0.6+0.6 +7.1+7.1
Fine-Tuning
vs ViT-MAE +1.2+1.2 1.9-1.9 +2.0+2.0
vs strongest DINOv2 +2.9+2.9 3.1-3.1 +1.3+1.3
Table 15 summarizes the delta comparisons for both transfer settings. Relative to ViT-MAE, the gains are positive on all three datasets under linear probing and remain positive on MiraBest and RGZ DR1 under fine-tuning. Relative to the strongest DINOv2 baseline, the gains are selective: positive on LoTSS DR2 and RGZ DR1 under probing, and on MiraBest and RGZ DR1 under fine-tuning. The smaller fine-tuning deltas indicate that supervised adaptation closes part of the representation gap observed under linear probing. Taken together with the targeted DINOv2 ablation, these results support the ViT-MAE-based two-stage checkpoint as the primary STRADAViT release: it remains competitive with the stronger DINOv2-based alternative while retaining a substantially lower token count and lower downstream cost. Figure 4 decomposes these aggregate results at class level.
Figure 4 compares the selected STRADAViT configuration against the two starting-point baselines, ViT-Mae and DINOv2 (Registers), using aggregate recall-form confusion matrices on all three evaluation datasets. Panel (a) reports linear probing and panel (b) reports full fine-tuning. This comparison makes the classwise effects of the selected model explicit and clarifies the classwise structure underlying the observed Macro-F1 gains.
Figure 4: Aggregate recall-form confusion matrices (%) for MiraBest, LoTSS DR2, and RGZ DR1. Panel (a) shows the linear-probe comparison and panel (b) shows the full fine-tuning comparison. In both panels, the selected STRADAViT configuration is compared against the two starting-point baselines, ViT-Mae and DINOv2 (Registers). Rows denote true classes and columns denote predicted classes.
Under linear probing, shown in Figure 4(a), the clearest classwise advantage appears on RGZ DR1: the selected model strengthens all six diagonal entries, reaching 92%, 81%, 78%, 85%, 72%, and 87% from 1C1P through 3C3P. The gain is therefore distributed across the full label space rather than concentrated in a single easy row. MiraBest is also favorable in this comparison, with the selected model reaching 68% on FR I and 72% on FR II, ahead of both ViT-Mae and DINOv2 (Registers). LoTSS DR2 remains mixed: the selected model leads on FR I, FR II, and Spiral, but the Hybrid and Relaxed-double rows remain strongest for ViT-Mae.
Under full fine-tuning, shown in Figure 4(b), the classwise margins are smaller. On MiraBest, all three models are tied at 70% on FR I, while the selected model matches the strongest FR II row at 75%. LoTSS DR2 remains the clearest limitation: the selected model is strongest only on FR II (88%) and trails at least one baseline on the remaining rows. RGZ DR1 remains partly favorable, but the gain is more localized than under probing: the selected model matches the best 1C1P row at 93% and leads on 1C2P and 2C2P (90% and 94%), whereas DINOv2 (Registers) remains stronger on 1C3P, 2C3P, and 3C3P. The fine-tuning deltas are therefore associated with selective classwise changes rather than with a uniform shift of the full matrix.
10.4 Overall Interpretation and Limitations
The branch comparison indicates that additional stages are not uniformly beneficial across settings. Reconstruction-only continued pretraining is the weakest of the three branches and often underperforms the ViT-MAE starting point. Masked reconstruction alone is therefore not sufficient to improve transfer in this setting. Introducing contrastive learning changes the picture substantially: contrastive-only training already produces large gains over phase 1 alone, especially under linear probing. Across the branch comparison, the two-stage branch is the most consistently top-performing STRADAViT variant, even though a small number of settings—most notably fine-tuned RGZ DR1—are matched or marginally exceeded by contrastive-only training.
The comparison against the actual initialization is informative in two respects. First, relative to ViT-MAE, the strongest two-stage models improve on all linear-probe settings and on fine-tuned MiraBest and RGZ DR1. Second, the branch comparison shows that most of this benefit comes from the contrastive objective, while the reconstruction stage contributes a smaller, dataset-dependent benefit. Phase 1 is best interpreted as a domain-adaptive warm start or regularizer for the subsequent contrastive stage, not as the main source of representation quality by itself. Two factors appear most influential: contrastive losses and ROI-aware view generation. The former better fits a setting in which many negatives are easy background cases while a smaller set are genuinely confusable; the latter makes both contrastive learning and masked reconstruction workable on sparse radio astronomy cutouts. This difference between frozen-feature gains and end-to-end gains is explicit in Table 15 and Figure 4: linear-probe improvements are larger and more consistent across datasets, whereas fine-tuning gains are smaller, more localized to specific classes, and change sign on LoTSS DR2. The classwise comparisons therefore indicate broader representation-level gains under probing than under end-to-end adaptation.
Across the contrastive objectives, InfoNCE provides a baseline but rarely defines the strongest probe results. Soft-HCL is the most consistent choice on LoTSS DR2 and in many register-based settings, whereas HCL tends to give the best MiraBest results and remains strongest on RGZ DR1 fine-tuning. Register-token effects are also more favorable inside the STRADAViT pipeline than in the off-the-shelf DINO baselines, especially on LoTSS DR2 and RGZ DR1 under probing, although the gain is not uniform across every branch and dataset. The DINOv2-Base(R)-initialized ablation is consistent with this pattern: it indicates portability of the contrastive adaptation recipe to a stronger starting point, but the resulting gains remain small and dataset-dependent, with the same LoTSS DR2 limitation that appears across the adapted models more generally.
The view generator is strongly object-centric. This property reduces empty crops and supports training on compact or moderately extended sources, but it also biases the pipeline toward RGZ-style views. LoTSS DR2 and MiraBest contain larger-scale extended structure that benefits from larger-view sampling designed for extended source morphology in radio astronomy.
11 Conclusions and Future Work
This paper studies how to adapt self-supervised ViT continued pretraining to sparse, heterogeneous radio astronomy imaging. The main result is that masked reconstruction alone is insufficient, contrastive learning is the main source of transfer gain, and the two-stage reconstruction-to-contrastive recipe is the most consistent STRADAViT configuration overall. The targeted DINOv2 ablation indicates that this adaptation recipe is not specific to the ViT-MAE starting point, but the ViT-MAE-based two-stage checkpoint remains the preferred STRADAViT release because it combines competitive transfer with substantially lower token count and lower downstream cost than the DINOv2-based alternative. In that sense, STRADAViT is best interpreted as a domain-adapted starting point for radio astronomy vision backbones: it provides a better practical replacement for out-of-the-box ViTs in frozen-backbone transfer, and remains favorable in most fine-tuning settings without implying uniform superiority in every dataset or class. We release this ViT-MAE-based checkpoint on Hugging Face [^1] https://huggingface.co/ISSA-ML/stradavit-base. Code, configuration files, and dataset-construction scripts are available from the corresponding author upon reasonable request; where direct redistribution of underlying data products is not possible, the reconstruction pipeline and source references can be provided.
Future work should test stronger teacher–student pretraining schemes, especially DINO-style variants built on anchored radio astronomy-aware view generation, together with larger-view regimes that preserve extended emission more consistently in pretraining. Further iterations of the on-the-fly data-curation and view-generation pipeline are also warranted, especially through augmentation and sampling schemes that better match the heterogeneity of the downstream datasets rather than remaining predominantly RGZ-like, while maintaining the general recipe presented here. Such refinements should improve the cross-dataset balance of the learned representations and are therefore a plausible route to more uniform downstream gains. The strong linear-probe gains also motivate evaluation in fixed-backbone settings, including detection and segmentation pipelines in which STRADAViT can serve as a domain-adapted feature extractor in place of weaker non-specialized ViT backbones. In parallel, the stable downstream gains observed under full fine-tuning justify further evaluation of task-specific models built directly on top of STRADAViT. Additional priorities are broader cross-survey robustness tests, explicit out-of-domain transfer, and repetition of the same controlled pipeline from stronger off-the-shelf initializations.
Acknowledgements
The authors acknowledge Xjenza Malta for funding this study under the Research Excellence Programme of 2024 [REP-2024-19]. The authors thank Alessio Magro and Kristian Grixti, Institute of Space Sciences and Astronomy, University of Malta, for technical support and for maintaining the GPU computing infrastructure used for the training and evaluation runs in this study.
[…]

### 08 Apr 2026 v3

@@
 Abstract
-Next-generation radio astronomy surveys are delivering millions of resolved sources, but scientific exploitation is increasingly limited by robust, scalable morphology analysis across heterogeneous telescopes and imaging pipelines.
+Next-generation radio astronomy surveys are delivering millions of resolved sources, but robust and scalable morphology analysis remains difficult across heterogeneous telescopes and imaging pipelines.
 We present STRADAViT, a self-supervised Vision Transformer (ViT) continued-pretraining framework for learning transferable encoders from radio astronomy imagery.
-The framework combines mixed-survey data curation, radio astronomy-aware training-view generation, and a ViT-MAE backbone with optional register tokens.
+The framework combines mixed-survey data curation, radio astronomy-aware training-view generation, and a ViT-MAE-initialized encoder family with optional register tokens.
 It supports reconstruction-only, contrastive-only, and two-stage branches.
-Our pretraining dataset comprises $512\times 512$ radio astronomy images drawn from four complementary sources (MeerKAT, ASKAP, LOFAR/LoTSS, and SKA).
+Our pretraining dataset comprises $512\times 512$ radio astronomy cutouts drawn from four complementary sources (MeerKAT, ASKAP, LOFAR/LoTSS, and SKA SDC1 simulated data).
 We evaluate transfer with linear probing and fine-tuning on three morphology benchmarks spanning binary and multi-class settings (MiraBest, LoTSS DR2, and Radio Galaxy Zoo).
 Relative to the ViT-MAE initialization used for continued pretraining, the best two-stage models improve Macro-F1 in all reported linear-probe settings and in two of three fine-tuning settings, with the largest gain on RGZ DR1.
-Relative to DINOv2, gains are selective rather than universal: the best two-stage models surpass the strongest DINOv2 baseline on LoTSS DR2 and RGZ DR1 under linear probing, and on MiraBest and RGZ DR1 under fine-tuning.
-A targeted DINOv2-initialized HCL ablation further indicates that the adaptation recipe is not specific to the ViT-MAE starting point.
-The ViT-MAE-based STRADAViT checkpoint nevertheless remains the preferred release because it combines competitive transfer with substantially lower token count and downstream cost than the DINOv2-based alternative.
-These results indicate that radio astronomy-aware view generation and staged continued pretraining can provide a stronger domain-adapted starting point than out-of-the-box ViT checkpoints for radio astronomy transfer, especially when representation quality is assessed through linear probing.
+Relative to DINOv2, gains are selective rather than universal: the best two-stage models achieve higher mean Macro-F1 than the strongest DINOv2 baseline on LoTSS DR2 and RGZ DR1 under linear probing, and on MiraBest and RGZ DR1 under fine-tuning.
+A targeted DINOv2 initialization ablation further indicates that the adaptation recipe is not specific to the ViT-MAE starting point and that, under the same HCL recipe, the register-based DINOv2 initialization is stronger than the non-register alternative.
+The ViT-MAE-based STRADAViT checkpoint is retained as the released checkpoint because it combines competitive transfer with substantially lower token count and downstream cost than the DINOv2-based alternative.
+These results indicate that radio astronomy-aware view generation and staged continued pretraining can provide a stronger domain-adapted starting point than off-the-shelf ViT checkpoints for radio astronomy transfer, especially when representation quality is assessed through linear probing.
 keywords:
 […]
-{highlights}
-Mixed-survey self-supervised dataset spanning MeerKAT, ASKAP, LoTSS, and SKA-derived cutouts
-Radio astronomy-aware view generation for sparse and heterogeneous radio astronomy images
-Controlled comparison of reconstruction-only, contrastive-only, and two-stage continued pretraining
-Two-stage STRADAViT improves on the ViT-MAE starting point in all reported linear-probe settings and in most fine-tuning settings
-Domain-adapted transfer is strongest under frozen-backbone evaluation, with selective gains against DINOv2 baselines
 1 Introduction
@@
 In mainstream computer vision, transferable representations are increasingly learned via self-supervised pretraining at scale and then adapted to diverse downstream tasks.
-Vision Transformers (ViTs) build on the Transformer paradigm (Vaswani et al., 2017) and have proven effective when scaled and pretrained appropriately (Dosovitskiy et al., 2021).
-Self-supervised objectives such as masked image modeling (He et al., 2022) and contrastive or invariance-based learning (Chen et al., 2020; He et al., 2020; Grill et al., 2020; Bardes et al., 2022) enable representation learning without labels, while recent work demonstrates strong ViT features learned purely from self-supervision (Caron et al., 2021; Oquab et al., 2023).
+Vision Transformers (ViTs) build on the Transformer paradigm (Vaswani2017Attention) and have proven effective when scaled and pretrained appropriately (Dosovitskiy2021ViT).
+Self-supervised objectives such as masked image modeling (He2022MAE) and contrastive or invariance-based learning (Chen2020SimCLR; He2020MoCo; Grill2020BYOL; Bardes2022VICReg) enable representation learning without labels, while recent work demonstrates strong ViT features learned purely from self-supervision (Caron2021DINO; Oquab2023DINOv2).
 A key open question for radio astronomy is how to adapt these pretraining ideas under the constraints of scientific imaging: single-channel data, survey-dependent intensity distributions, and heterogeneous instrumental/imaging systematics.
-This paper presents STRADAViT, developed as part of STRADA (Self-supervised Transformers for Radio Astronomy Discovery Algorithms).
-The framework is organized around three components: (i) a mixed-survey self-supervised pretraining dataset drawn from four complementary radio astronomy image sources, (ii) radio astronomy-aware training-view generation for masked reconstruction and contrastive learning, and (iii) a continued-pretraining framework built on ViT-MAE.
-The framework supports reconstruction-only, contrastive-only, and two-stage branches.
+This paper presents STRADAViT, a self-supervised transfer framework for radio astronomy Vision Transformers.
 Our primary contributions are:
@@
   • radio astronomy-aware training-view generation that anchors self-supervised views to informative source regions instead of relying on naive random crops;
-  • a ViT-MAE-based self-supervised framework with controlled ablations of reconstruction losses, hard-negative-aware contrastive losses, and register tokens, evaluated through reconstruction-only, contrastive-only, and two-stage continued pretraining;
+  • a controlled continued-pretraining framework initialized from ViT-MAE, with optional register tokens and controlled ablations of reconstruction losses, hard-negative-aware contrastive losses, and reconstruction-only, contrastive-only, and two-stage training branches;
+  • a targeted DINOv2 initialization ablation under a fixed contrastive recipe, used to test portability of the adaptation pipeline across materially different ViT starting points;
   • a reproducible transfer protocol (linear probing and fine-tuning) across three public morphology benchmarks, with analysis emphasizing the probe–fine-tune gap and dataset sensitivity.
@@
 Sections 3 and 4 describe the pretraining and evaluation datasets.
-Sections 57 define the model, training branches, and pretraining objectives.
-Transfer results are then presented under linear probing and fine-tuning.
-We then present classwise confusion-pattern comparisons against the starting-point baselines, followed by an overall interpretation of the branch comparison and the conclusion.
+Sections 57 define the model, training branches, and pretraining objectives, and Section 8 summarizes initialization, optimization, and branch-specific training settings.
+Section 9 then defines the downstream transfer protocol for linear probing and fine-tuning.
+Section 10 presents the baseline comparisons, continued-pretraining results, the targeted DINOv2 initialization ablation, and the classwise comparison for the selected release checkpoint, followed by the overall interpretation and the conclusion.
 2 Literature Survey
-This survey covers the transferable-representation literature most directly relevant to the present design choices.
-2.1 Backbones for Transferable Visual Representations
-Convolutional neural networks remain strong imaging baselines, particularly in residual form (He et al., 2016), but Transformer-based backbones now dominate large-scale representation learning (Vaswani et al., 2017; Dosovitskiy et al., 2021).
-In specialized imaging domains, backbone choice interacts closely with pretraining objective, dataset composition, and normalization.
-2.2 Self-Supervised Learning: Contrastive Invariances and Masked Modeling
+Convolutional neural networks remain strong imaging baselines, particularly in residual form (He2016ResNet), but Transformer-based backbones now dominate large-scale representation learning (Vaswani2017Attention; Dosovitskiy2021ViT).
+In specialized imaging domains, backbone choice interacts closely with pretraining objective, dataset composition, and normalization, which motivates keeping the backbone family controlled so that transfer differences can be attributed primarily to the pretraining recipe and view generation.
 Self-supervised learning replaces manual labels with surrogate objectives.
-Contrastive methods such as SimCLR and MoCo learn invariances across augmented views (Chen et al., 2020; He et al., 2020), whereas related methods modify the handling of negatives or regularize feature statistics (Grill et al., 2020; Bardes et al., 2022).
-Masked image modeling provides a complementary signal: MAE reconstructs masked content from partial observations and is effective when image statistics differ substantially from those of natural images (He et al., 2022).
-DINO-style methods and DINOv2 further show that training recipe and architectural details, including register tokens, materially affect transfer (Caron et al., 2021; Oquab et al., 2023).
-These two objective families motivate the staged design studied here.
-2.3 From General-Purpose Pretraining to Domain-Specific Transfer
-General-purpose pretraining can yield strong transferable features (Radford et al., 2021), but radio astronomy differs markedly from natural-image domains: images are typically single-channel, have survey-dependent noise and dynamic range, and contain instrument- and pipeline-specific artefacts.
+Contrastive methods such as SimCLR and MoCo learn invariances across augmented views (Chen2020SimCLR; He2020MoCo), whereas related methods modify the handling of negatives or regularize feature statistics (Grill2020BYOL; Bardes2022VICReg).
+Masked image modeling provides a complementary signal: MAE reconstructs masked content from partial observations and is effective when image statistics differ substantially from those of natural images (He2022MAE).
+DINO-style methods and DINOv2 further show that training recipe and architectural details, including register tokens, materially affect transfer (Caron2021DINO; Oquab2023DINOv2).
+These two objective families motivate the staged reconstruction-to-contrastive design studied here.
+General-purpose pretraining can yield strong transferable features (Radford2021CLIP), but radio astronomy differs markedly from natural-image domains: images are typically single-channel, have survey-dependent noise and dynamic range, and contain instrument- and pipeline-specific artefacts.
 Effective transfer therefore depends on representative unlabeled data, domain-appropriate normalization, and augmentations that preserve morphology rather than photographic appearance.
-2.4 Radio Astronomy: From Source Finding to Transferable Encoders
-Beyond morphology classification, radio astronomy has a long tradition of source finding and characterization under survey-specific artefacts and systematics, as illustrated by CAESAR and related practical pipelines (Riggi et al., 2016, 2019).
-Recent work extends this toward broad reviews of deep learning for radio classification (Riggi et al., 2024), self-supervised transfer and benchmark studies (Riggi et al., 2024; Cecconello et al., 2025), and multimodal systems (Riggi et al., 2025; Drozdova et al., 2025).
-Additional SSL studies now cover source classification, FR I/FR II transfer, foundation-style pretraining on Radio Galaxy Zoo, and self-supervised learning on MeerKAT continuum images (Baron Pérez et al., 2025; Buatthaisong et al., 2025; Slijepcevic et al., 2023; Lastufka et al., 2024).
-2.5 Radio Morphology Datasets and Benchmark Labels
+2.1 Radio Astronomy Context and Benchmark Labels
+Beyond morphology classification, radio astronomy has a long tradition of source finding and characterization under survey-specific artefacts and systematics, as illustrated by CAESAR and related practical pipelines (Riggi2016CAESAR; Riggi2019CaesarPASA).
+Recent work extends this toward broad reviews of deep learning for radio classification (Riggi2024DeepLearningRadio), self-supervised transfer and benchmark studies (Riggi2024PASAContrastive; Cecconello2025RadioSSLBenchmark), and multimodal systems (Riggi2025RadioLLaVA; Drozdova2025VLMRadio).
+Additional SSL studies now cover source classification, FR I/FR II transfer, foundation-style pretraining on Radio Galaxy Zoo, and self-supervised learning on MeerKAT continuum images (BaronPerez2025AandASSL; Buatthaisong2025MNRASSelfSupervisedFR; Slijepcevic2024RGZFoundation; Lastufka2024MeerKATSSL).
 Radio morphology benchmarks remain anchored in expert or citizen-science labels.
-LoTSS DR2 provides broad morphology labels and feature flags for bright, extended sources (Horton et al., 2025); MiraBest provides curated FR I/FR II splits (Porter and Scaife, 2023); and Radio Galaxy Zoo operationalizes large-scale citizen-science labeling for more complex multi-class settings (Wong et al., 2024).
+LoTSS DR2 provides broad morphology labels and feature flags for bright, extended sources (Horton2025LoTSSDR2Morphology); MiraBest provides curated FR I/FR II splits (PorterScaife2023MiraBest); and Radio Galaxy Zoo operationalizes large-scale citizen-science labeling for more complex multi-class settings (RGZDR1Zenodo14195049).
 Across these datasets, labels cover only part of the available data and label spaces are not harmonized across surveys, which makes cross-survey evaluation essential.
 2.2 Positioning Relative to Recent Radio Astronomy SSL Literature
-Relative to recent radio astronomy SSL studies (Riggi et al., 2024; Cecconello et al., 2025; Lastufka et al., 2024; Baron Pérez et al., 2025; Buatthaisong et al., 2025; Slijepcevic et al., 2023), STRADAViT emphasizes four points: mixed-survey pretraining, ROI-aware view generation for sparse cutouts, a staged reconstruction-to-contrastive objective, and a controlled ViT-only architectural scope.
+Relative to recent radio astronomy SSL studies (Riggi2024PASAContrastive; Cecconello2025RadioSSLBenchmark; Lastufka2024MeerKATSSL; BaronPerez2025AandASSL; Buatthaisong2025MNRASSelfSupervisedFR; Slijepcevic2024RGZFoundation), STRADAViT emphasizes four points: mixed-survey pretraining, ROI-aware view generation for sparse cutouts, a staged reconstruction-to-contrastive objective, and a controlled ViT-only architectural scope.
 The aim is cross-survey reuse across multiple downstream label spaces rather than optimization for a single benchmark.
@@
 Our self-supervised pretraining dataset is constructed to reflect the heterogeneity of contemporary radio astronomy imaging across instruments and imaging pipelines.
-We assemble $512\times 512$ continuum-image cutouts from four complementary sources: MeerKAT MGCLS DR1, ASKAP continuum images accessed via the CSIRO ASKAP Science Data Archive, LOFAR LoTSS continuum mosaics, and simulated continuum images from the first SKA Science Data Challenge (SDC1) DR1 (Bonaldi et al., 2020).
+We assemble $512\times 512$ continuum-image cutouts from four complementary sources: MeerKAT MGCLS DR1, ASKAP continuum images accessed via the CSIRO ASKAP Science Data Archive, LOFAR LoTSS continuum mosaics, and simulated continuum images from the first SKA Science Data Challenge (SDC1) DR1 (Bonaldi2020).
+These sources cover different angular resolutions, noise properties, imaging artefacts, and source morphologies, so the pretraining dataset is intended to expose the model to heterogeneous radio structure rather than to a single survey regime.
+Table 1 summarizes the resulting dataset size after preprocessing and filtering.
 3.1 Cutout Extraction and Basic Filtering
@@
 For compatibility with standard vision backbones, the resulting single-channel image is replicated to three channels after scaling.
-3.3 Dataset Composition
-Table 1 summarizes the resulting dataset size after preprocessing and filtering.
 Table 1: Composition of the self-supervised pretraining dataset (all images are $512\times 512$ cutouts). “Ratio” denotes the fraction of total cutouts contributed by each source.
@@
 In the contrastive loss, the two views of the same cutout form the positives, while all other views in the (distributed) micro-batch act as negatives.
-3.4 Online View Generation and Extensibility
+3.3 Online View Generation
 All views used for self-supervision are generated on-the-fly during training (Sections 6 and 7), rather than being precomputed.
-As a consequence, a given cutout is typically encountered under many different ROI selections, crop scales, and augmentations over the course of training, which increases effective data diversity without requiring additional offline curation.
-This runtime design is also intended to make the pretraining dataset extensible: additional telescope or simulation sources can be incorporated by adding new mosaics (or cutouts) that satisfy the same basic FITS tiling and preprocessing interface, without redesigning survey-specific curation rules.
+A given cutout is therefore typically encountered under many different ROI selections, crop scales, and augmentations over the course of training, which increases effective data diversity without additional offline curation.
+The same runtime design also makes the pretraining dataset extensible, since additional telescope or simulation sources can be incorporated by adding new mosaics (or cutouts) that satisfy the same FITS tiling and preprocessing interface.
 4 Evaluation Datasets
-We evaluate transfer learning on three public, morphology-focused radio astronomy imaging datasets that span different telescopes, labeling paradigms, and class granularities: MiraBest (Porter and Scaife, 2023), the LoTSS DR2 visual-classification sample of Horton et al.
-(2025), and Radio Galaxy Zoo DR1 (Wong et al., 2024).
-Together, these datasets allow us to probe binary and multi-class settings and to test robustness across heterogeneous imaging products.
+We evaluate transfer learning on three public radio astronomy imaging datasets that span different telescopes, labeling paradigms, and class granularities: MiraBest (PorterScaife2023MiraBest), the LoTSS DR2 visual-classification sample of Horton2025LoTSSDR2Morphology, and Radio Galaxy Zoo DR1 (RGZDR1Zenodo14195049).
+Together, these datasets probe compact binary morphology, richer multi-class morphology, and component/peak complexity under heterogeneous imaging products.
 Table 2 summarizes the number of retained samples in each evaluation dataset after applying the dataset selection and filtering criteria described below.
@@
 Figure 3: Example cutouts from the three evaluation datasets after preprocessing with the same pipeline used for pretraining and evaluation (Section 3).
-Panels show samples of each class for (a) RGZ DR1, (b) MiraBest, and (c) the LoTSS DR2 visual-classification sample of Horton et al.
-(2025).
+Panels show samples of each class for (a) RGZ DR1, (b) MiraBest, and (c) the LoTSS DR2 visual-classification sample of Horton2025LoTSSDR2Morphology.
 4.1 Common Preprocessing
@@
 In downstream evaluation this resolution is model-dependent: ViT-MAE-family models use $224\times 224$ inputs, whereas DINOv2-family models use their native $518\times 518$ inputs.
-For FITS-based datasets, unreadable files and blank cutouts (e.g., all-zero images) are excluded during indexing.
 4.2 MiraBest (Binary FR I/FR II)
-MiraBest provides a curated set of morphologically classified radio galaxies intended for machine learning applications (Porter and Scaife, 2023).
+MiraBest provides a curated set of morphologically classified radio galaxies intended for machine learning applications (PorterScaife2023MiraBest).
 The released labels include multiple subtypes and confidence levels.
@@
 Unless stated otherwise, we further restrict to confident labels and discard uncertain entries to avoid conflating representation quality with label ambiguity.
-The images are distributed in a batched format; we apply the same normalization and resizing pipeline as for FITS cutouts to obtain a standardized tensor input.
-4.3 LoTSS DR2 Visual-Classification Cutouts (Multi-Class Initial Labels)
-Horton et al.
-(2025) present a large visual-classification effort for bright, resolved LoTSS DR2 sources, including an “initial” class taxonomy and additional morphology and environment flags.
+We apply the same normalization and resizing pipeline as for FITS cutouts to obtain a standardized tensor input.
+4.3 LoTSS DR2 (Multi-Class Initial Labels)
+Horton2025LoTSSDR2Morphology present a large visual-classification effort for bright, resolved LoTSS DR2 sources, including an “initial” class taxonomy and additional morphology and environment flags.
 We download high-resolution LoTSS DR2 cutouts for each source using the public cutout service, centring on the catalog coordinates.
@@
 Samples with no initial label (or ambiguous multiple initial labels) are excluded.
-We retain the full set of per-source flags in a JSON sidecar for reproducibility and for possible multi-label extensions.
-4.4 Radio Galaxy Zoo DR1 (Six-Class Component/Peak Morphology)
-Radio Galaxy Zoo DR1 provides citizen-science classifications for complex radio sources, including a consensus level that can be used to filter for high-confidence labels (Wong et al., 2024).
+4.4 Radio Galaxy Zoo DR1 (Multi-Class Component/Peak Morphology)
+Radio Galaxy Zoo DR1 provides citizen-science classifications for complex radio sources, including a consensus level that can be used to filter for high-confidence labels (RGZDR1Zenodo14195049).
 From the FIRST subset, we construct a multi-class benchmark based on the discrete component/peak summary provided in the catalog.
 Specifically, we derive a class label from the number of radio components and peaks and retain the six most common combinations: 1c1p, 1c2p, 1c3p, 2c2p, 2c3p, and 3c3p.
-We filter the catalog to consensus level $\mathrm{CL}\geq 0.65$ and download FIRST FITS cutouts via the FIRST cutout service, using an adaptive angular size of 1.5 times the catalog largest angular extent (LAE) and enforcing a minimum cutout side length of 8 pixels.
+We filter the catalog to consensus level $\geq 0.65$ and download cutouts via the FIRST cutout service, using an adaptive angular size of 1.5 times the catalog largest angular extent (LAE) and enforcing a minimum cutout side length of 8 pixels.
 To improve data integrity, cutouts are rejected when the WCS-derived image centre deviates from the target coordinate by more than 1 arcmin.
@@
 5.1 Base Encoder: ViT-MAE
-STRADAViT builds on the ViT-MAE formulation (He et al., 2022), which uses a Vision Transformer (ViT) encoder (Dosovitskiy et al., 2021; Vaswani et al., 2017) with a lightweight decoder during masked reconstruction.
-Images are split into non-overlapping $p\times p$ patches and linearly embedded into a sequence of token vectors.
-A learnable class token is prepended and positional encodings are added, after which the sequence is processed by a stack of Transformer encoder blocks (multi-head self-attention and a feed-forward MLP).
+STRADAViT builds on the ViT-MAE formulation (He2022MAE), which uses a Vision Transformer (ViT) encoder (Dosovitskiy2021ViT; Vaswani2017Attention) with a lightweight decoder during masked reconstruction.
+Images are split into non-overlapping $p\times p$ patches and linearly embedded into a sequence of token vectors, which is then processed by a stack of Transformer encoder blocks (multi-head self-attention and a feed-forward MLP).
 Unless stated otherwise, we use the standard ViT-MAE base configuration (ViT-B/16).
@@
 We choose this MAE-family starting point because it provides a single, readily modifiable encoder family for both reconstruction and contrastive branches, allowing a controlled study of the transfer-learning training recipe without introducing additional implementation and compute variance from a second pretrained stack.
-All training views are resized to the model input resolution ($224\times 224$).
 In the reconstruction objective, the decoder is used to reconstruct masked patches from the encoder representation.
@@
 5.3 Register Tokens
-We also evaluate an optional register-token variant inspired by recent foundation-style ViT designs (Oquab et al., 2023).
+We also evaluate an optional register-token variant inspired by recent foundation-style ViT designs (Oquab2023DINOv2).
 In this setting, we insert $R$ learnable register tokens immediately after the class token and before the patch tokens.
@@
 A square ROI is then sampled around this anchor with side length fraction in $[0.55,0.80]$ of the cutout size, and a “global-tight” view is sampled inside the ROI with scale range $[0.50,0.70]$ (resized to the model input size), with bounded resampling if an uninformative crop is proposed.
-After the geometric step, we apply the mild augmentation regime in Table 3, clamp to $[0,1]$, and then apply the initialization-dependent channel normalization described in Section 8.1.
+After the geometric step, we apply the mild augmentation regime in Table 3, clamp to $[0,1]$, and then apply model-dependent channel normalization (i.e., ImageNet normalization for pretrained backbones).
 Figure 1 illustrates representative reconstruction-branch views produced by this pipeline.
@@
 6.2 Reconstruction Loss Variants
-The reconstruction branch follows the ViT-MAE masked image modeling objective (He et al., 2022).
+The reconstruction branch follows the ViT-MAE masked image modeling objective (He2022MAE).
 Let $x$ denote an input view, $m$ a binary mask over patch locations, and $\hat{x}$ the model reconstruction (after unpatchifying decoder outputs).
@@
 $\mathcal{L}_{\mathrm{MAE}}(x,\hat{x};m)=\frac{1}{|m|}\sum_{j\in m}\lVert\hat{x}_{j}-x_{j}\rVert_{2}^{2}.$ (2)
-We treat this as our baseline and compare two explicit modifications that add an additional reconstruction regularizer computed in the same normalized input space as $x$.
+We treat this as our baseline, denote it by $\mathcal{L}_{\mathrm{L2}}$, and set its weight to $w_{\mathrm{L2}}=1.0$:
+$\mathcal{L}_{\mathrm{L2}}=w_{\mathrm{L2}}\,\mathcal{L}_{\mathrm{MAE}},$ (3)
+We then compare two explicit modifications that add an additional reconstruction regularizer computed in the same normalized input space as $x$.
 In the following, $\langle\cdot\rangle$ denotes the mean over all pixels (and channels) of an image.
-6.2.1 Baseline: Masked-Patch MSE (ViT-MAE Default)
-The baseline objective is
-$\mathcal{L}_{\mathrm{L2}}=w_{\mathrm{L2}}\,\mathcal{L}_{\mathrm{MAE}},$ (3)
-with weight $w_{\mathrm{L2}}=1.0$.
 6.2.1 Masked-Patch MSE + Global L1
@@
 We apply the contrastive objective to the $\ell_{2}$-normalized projection-head outputs (Section 5), using all other views in the (distributed) batch as negatives.
-We compare three contrastive loss variants: a standard InfoNCE formulation, a “soft” hard-negative reweighting variant (soft-HCL), and a hard-negative objective (HCL) (Robinson et al., 2021).
+We compare three contrastive loss variants: a standard InfoNCE formulation, a “soft” hard-negative reweighting variant (soft-HCL), and a hard-negative objective (HCL) (Robinson2021HardNegatives).
 7.1 View Generation for Contrastive Learning
@@
 Augmentations follow the same family as in the reconstruction branch, with the strong regime applied to the additional view.
-Each view is clamped to $[0,1]$ and normalized using the policy in Section 8.1.
+Each view is clamped to $[0,1]$ and normalized using the same model-dependent policy.
 7.1.1 Online Non-Empty Resampling
@@
 Each training step operates on a batch of $B$ cutouts, generating $V$ views per cutout (in this work $V=2$) and producing an embedding $z\in\mathbb{R}^{d}$ for each view via the encoder, mean pooling over patch tokens, and the projection head (Section 5).
-Embeddings are $\ell_{2}$-normalized and concatenated across all devices to form a global set of keys; for each query embedding, positives are the other view(s) of the same cutout and negatives are all remaining keys (except the query itself).
 Similarities are computed as dot products and scaled by a temperature $\tau$ that is linearly warmed from $0.20$ to $0.10$ over the warmup portion of training (first 15% of epochs) and then kept at $\tau=0.10$.
-In our distributed implementation, embeddings are all-gathered across devices to form the key set, but embeddings from other devices are treated as stop-gradient keys (i.e., gathered without gradient) while the loss is evaluated for the local queries.
-The size of the negative pool is bounded by the distributed micro-batch used to compute the loss (i.e., a single forward pass).
-Let $G$ denote the number of devices (world size).
-Each loss evaluation involves $G\!\cdot\!B\!\cdot\!V$ key embeddings (all views across devices).
-For a given query, the number of negative keys is therefore approximately $G\!\cdot\!B\!\cdot\!V-V$ (excluding the $V$ views of the query’s underlying cutout, including itself).
-In the runs reported in this work we use $G=4$ GPUs, $B=64$ images per GPU, and $V=2$ views, yielding 256 images (512 keys) per micro-batch and 510 negative keys per query (equivalently, 255 other images).
-We use gradient accumulation to obtain a fixed effective global batch size of 2048 images per optimizer update.
-However, gradient accumulation does not increase the negative pool per loss evaluation.
-InfoNCE, HCL and soft-HCL share the same negative pool; they differ only in how negatives are weighted.
+In our distributed implementation, the normalized embeddings are all-gathered across devices to form the key set, but embeddings from other devices are treated as stop-gradient keys (i.e., gathered without gradient) while the loss is evaluated for the local queries.
+In the ViT-MAE-based runs reported here, 4 GPUs with per-device batch size 64 and 2 views per cutout yield 512 keys per micro-batch and 510 negatives per query; gradient accumulation increases the effective optimizer batch size but does not enlarge this per-step negative pool.
+InfoNCE, HCL, and soft-HCL therefore share the same negative pool and differ only in how negatives are weighted.
 7.2.1 Baseline: InfoNCE
-We use a standard temperature-scaled InfoNCE contrastive loss (the NT-Xent form used in SimCLR (Chen et al., 2020)) as our baseline.
+We use a standard temperature-scaled InfoNCE contrastive loss (the NT-Xent form used in SimCLR (Chen2020SimCLR)) as our baseline.
 Let $s_{ij}=z_{i}^{\top}z_{j}$ denote the similarity between a query $z_{i}$ and a key $z_{j}$.
@@
 7.2.2 Hard Contrastive Learning (HCL)
-We also evaluate hard-negative reweighting following Robinson et al.
-(2021).
+We also evaluate hard-negative reweighting following Robinson2021HardNegatives.
 HCL keeps the same overall contrastive log-ratio structure as InfoNCE, but it modifies how the negative term is accumulated: negatives that are more similar to the query contribute more to the denominator.
@@
 8 Training Setup
-This section summarizes (i) the initialization and input-normalization choices and (ii) the optimization settings and training schedules used in the reconstruction and contrastive branches.
-8.1 Initialization and Input Normalization
-We consider two starting points for continued pretraining: (i) continued pretraining from publicly available ViT-MAE weights pretrained on ImageNet-1K, and (ii) scratch training with random initialization.
-The choice of starting point also determines whether we apply an additional channel-wise input normalization step.
-When continuing from ImageNet-pretrained weights, we apply the standard ImageNet mean and standard deviation normalization to match the pretrained backbone.
-When training from scratch, we omit this normalization (identity), since our preprocessing already maps inputs to $[0,1]$.
+This section summarizes the optimization settings and training schedules used in the reconstruction and contrastive branches.
+All runs use continued pretraining from publicly available ViT-MAE weights pretrained on ImageNet-1K, with standard ImageNet mean and standard deviation normalization applied to match the pretrained backbone.
 The contrastive-only branch initializes directly from the chosen starting checkpoint, whereas the two-stage branch initializes its contrastive stage from the corresponding reconstruction checkpoint trained with the same architectural settings.
-8.2 Optimization and Training Hyperparameters
 Unless stated otherwise, settings are shared across runs and the effective global batch size is fixed at 2048 images per optimizer update via gradient accumulation.
-8.2.1 Common Optimization Settings
+8.1 Common Optimization Settings
 […]
-8.2.2 Distributed Training and Normalization Details
+8.2 Distributed Training and Normalization Details
 Training is performed using data-parallel distributed training across GPUs.
@@
 Because the projection head includes a BatchNorm layer, we use SyncBatchNorm across the 4 GPUs to ensure consistent normalization statistics.
-8.2.3 Reconstruction-Branch Training
-Unless stated otherwise, the reconstruction branch trains for 35 epochs with a fixed mask ratio $0.75$ and patch size $16$.
+8.3 Reconstruction-Branch Training
+The reconstruction branch trains for 35 epochs with a fixed mask ratio $0.75$ and patch size $16$.
 We use per-device batch size 256 and apply gradient clipping with $\lVert g\rVert_{2}$ capped at 1.0.
 We compare the three reconstruction losses described in Section 6 (masked-patch MSE; MSE + L1; MSE + brightness-weighted L1), using the weights specified there.
-8.2.4 Contrastive-Branch Training
-Unless stated otherwise, the contrastive branch trains for 35 epochs, both when used directly in the contrastive-only branch and as the second stage of the two-stage branch.
+8.4 Contrastive-Branch Training
+The contrastive branch trains for 35 epochs, both when used directly in the contrastive-only branch and as the second stage of the two-stage branch.
 We generate $V=2$ views per cutout and use per-device batch size 64 (so each step processes $2B$ views).
@@
   3. 3.
-    Input normalization: Channel-wise normalization uses the mean and standard deviation stored with each checkpoint’s image processor: ImageNet mean/std for ImageNet-initialized checkpoints and identity normalization for scratch-trained checkpoints (Section 8.1).
+    Input normalization: Channel-wise normalization uses the mean and standard deviation stored with each checkpoint’s image processor, i.e. ImageNet mean/std for ImageNet-initialized checkpoints.
   4. 4.
@@
 We therefore interpret probe gains as evidence of improved linear separability of the learned representations—a desirable property for transfer-oriented reuse—while treating fine-tuning as the primary measure of practical downstream performance on these benchmarks.
-All baselines and continued-pretraining variants are evaluated with the same downstream preprocessing, folds, and optimization protocol so that these comparisons remain controlled.
+In other words, linear probing serves as the sharper representation-quality diagnostic, whereas fine-tuning serves as the practical downstream check.
 10.1 Baselines (DINOv2, ViT-MAE)
@@
 By contrast, the register-token DINOv2 variant is weaker on every benchmark in the frozen-feature setting, with the largest drop on RGZ DR1 ($0.661\pm 0.004\rightarrow 0.629\pm 0.001$).
-The strongest off-the-shelf starting point in terms of raw linear separability is therefore the non-register DINOv2 backbone.
+The strongest off-the-shelf baseline under linear probing is therefore the non-register DINOv2 backbone.
 Fine-tuning yields a less uniform ranking (Table 7).
@@
 On LoTSS DR2, DINOv2-Base is retained over DINOv2-Base(R) after a Macro-F1 tie at $0.708$ because it has the higher Weighted-F1.
-The register variant therefore no longer dominates the fine-tuning baselines, although it remains the strongest DINOv2 model on RGZ DR1.
 Optimization sensitivity also remains pronounced: the non-register DINOv2-Base variant is unstable on MiraBest ($0.461\pm 0.195$), whereas DINOv2-Small ($0.707\pm 0.031$), DINOv2-Base(R) ($0.704\pm 0.102$), and ViT-MAE-Base ($0.724\pm 0.002$) are more stable there.
@@
 10.2.1 Reconstruction-Only Branch
-In this section we evaluate continued pretraining runs initialized from ImageNet-pretrained ViT-MAE weights (Section 8.1).
+In this section we evaluate continued pretraining runs initialized from ImageNet-pretrained ViT-MAE weights.
 These phase 1 results compare the reconstruction losses from Section 6 while keeping the rest of the protocol fixed.
@@
 We compare InfoNCE, Soft-HCL, and HCL under the same downstream protocol for the ViT-MAE-based runs.
-The separate DINOv2 initialization study is reported below under a fixed HCL recipe.
 Tables 10 and 11 summarize the contrastive-only results for the ViT-MAE-based runs, stratified by the presence of register tokens ($R\in\{0,4\}$).
@@
 HCL $0.722\pm 0.013$ $0.723\pm 0.013$ $0.665\pm 0.003$ $0.797\pm 0.003$ $0.826\pm 0.001$ $0.924\pm 0.001$
-The contrastive branch improves substantially on the phase 1 branch under both linear probing and fine-tuning.
+The contrastive branch yields substantial improvements over phase 1 under both linear probing and fine-tuning.
 The main gain therefore comes from the contrastive objective rather than from masked reconstruction.
 Loss rankings remain dataset-dependent: Soft-HCL is strongest in several MiraBest and LoTSS DR2 settings, whereas HCL is the most reliable choice on RGZ DR1.
-Registers interact with the loss in the same dataset-dependent way and do not produce a uniform ordering across datasets.
+This pattern is consistent with differences in class structure and confusability across datasets, and registers interact with the loss in the same dataset-dependent way rather than producing a uniform ordering.
 10.2.3 DINOv2 Initialization Ablation
 To test portability of the same adaptation pipeline to a differently pretrained ViT family, we also perform a targeted DINOv2 initialization ablation based on fixed HCL training.
-This provides a direct check of whether the radio astronomy adaptation pipeline transfers to a backbone whose original pretraining is already contrastive and more structured than ViT-MAE.
+This assesses whether the radio astronomy adaptation pipeline transfers to a backbone whose original pretraining is already contrastive and more structured than ViT-MAE.
 Table 12 summarizes this ablation.
 Table 12: DINOv2 initialization ablation under the contrastive-only HCL recipe (mean $\pm$ std over $K=3$ folds).
+Best Macro-F1 per dataset within each transfer setting is highlighted in bold; ties are broken by Weighted-F1.
 Setting Initialization MiraBest LoTSS DR2 RGZ DR1
 Macro-F1 Weighted-F1 Macro-F1 Weighted-F1 Macro-F1 Weighted-F1
-Linear Probe DINOv2-Base
+Linear Probe DINOv2-Base $0.683\pm 0.010$ $0.683\pm 0.010$ $0.551\pm 0.010$ $0.724\pm 0.006$ $0.735\pm 0.004$ $0.894\pm 0.000$
 Linear Probe DINOv2-Base(R) $0.686\pm 0.018$ $0.687\pm 0.018$ $0.561\pm 0.005$ $0.731\pm 0.003$ $0.739\pm 0.004$ $0.895\pm 0.000$
-Fine-Tuning DINOv2-Base
+Fine-Tuning DINOv2-Base $0.721\pm 0.020$ $0.722\pm 0.020$ $0.676\pm 0.009$ $0.816\pm 0.004$ $0.816\pm 0.000$ $0.919\pm 0.000$
 Fine-Tuning DINOv2-Base(R) $0.738\pm 0.014$ $0.738\pm 0.014$ $0.678\pm 0.015$ $0.816\pm 0.006$ $0.828\pm 0.002$ $0.922\pm 0.002$
-The DINOv2-based results reported here show that the same contrastive adaptation recipe is not specific to the ViT-MAE starting point.
-Under linear probing, the DINOv2-Base(R)-initialized run improves on the corresponding off-the-shelf DINOv2-Base(R) baseline on all three datasets, with the largest change on RGZ DR1.
-Under full fine-tuning, the adapted model improves on the off-the-shelf DINOv2-Base(R) baseline on MiraBest and RGZ DR1, but not on LoTSS DR2.
-Relative to the ViT-MAE-based adapted models, the DINOv2-initialized run is broadly competitive, with only small differences across all three datasets.
-The main exception remains LoTSS DR2, where the adapted DINOv2 result follows the same pattern as the ViT-MAE-based adapted models and remains below the off-the-shelf baselines.
+The DINOv2-based results show that the same contrastive adaptation recipe is not specific to the ViT-MAE starting point.
+These results should be read as a constrained initialization study, because the DINOv2 runs use a smaller micro-batch and therefore a smaller negative pool than the ViT-MAE runs.
+Under the fixed HCL recipe, DINOv2-Base(R) is the stronger adapted initialization on all three datasets in both transfer settings, even though the off-the-shelf linear-probe baselines favored DINOv2-Base without registers.
+The adaptation effect therefore depends on initialization family and does not simply preserve the off-the-shelf ranking.
+Relative to their own off-the-shelf baselines, the adapted DINOv2 variants improve on MiraBest and RGZ DR1, whereas LoTSS DR2 remains below the off-the-shelf DINOv2 results.
+Relative to the ViT-MAE-based adapted models, the best DINOv2-initialized run is broadly competitive and marginally stronger than the ViT-MAE-based contrastive-only models on MiraBest and LoTSS DR2 under fine-tuning.
+The ablation therefore supports portability of the adaptation recipe.
 10.2.4 Two-Stage Branch
@@
 Including the contrastive-only results clarifies the role of the reconstruction stage.
-Under linear probing, the best two-stage models remain the strongest STRADAViT configurations overall, but the extra reconstruction stage helps mainly on MiraBest and LoTSS DR2; RGZ DR1 is already well served by contrastive-only training.
+Under linear probing, the best two-stage models remain the strongest STRADAViT configurations overall and are therefore the basis of the selected release checkpoint, but the extra reconstruction stage helps mainly on MiraBest and LoTSS DR2; RGZ DR1 is already well served by contrastive-only training.
 Relative to the strongest off-the-shelf DINOv2 baseline, the best two-stage probe result remains below DINOv2 on MiraBest but exceeds it on LoTSS DR2 and RGZ DR1.
@@
 Relative to the strongest DINOv2 baseline, the gains are selective: positive on LoTSS DR2 and RGZ DR1 under probing, and on MiraBest and RGZ DR1 under fine-tuning.
-The smaller fine-tuning deltas indicate that supervised adaptation closes part of the representation gap observed under linear probing.
-Taken together with the targeted DINOv2 ablation, these results support the ViT-MAE-based two-stage checkpoint as the primary STRADAViT release: it remains competitive with the stronger DINOv2-based alternative while retaining a substantially lower token count and lower downstream cost.
-Figure 4 decomposes these aggregate results at class level.
+Together with the targeted DINOv2 ablation, these results support selection of the ViT-MAE-based two-stage checkpoint as the primary STRADAViT release: it remains competitive with the stronger DINOv2-based alternative while retaining a substantially lower token count and lower downstream cost.
 Figure 4 compares the selected STRADAViT configuration against the two starting-point baselines, ViT-MAE and DINOv2 (Registers), using aggregate recall-form confusion matrices on all three evaluation datasets.
@@
 Rows denote true classes and columns denote predicted classes.
-Under linear probing, shown in Figure 4(a), the clearest classwise advantage appears on RGZ DR1: the selected model strengthens all six diagonal entries, reaching 92%, 81%, 78%, 85%, 72%, and 87% from 1C1P through 3C3P.
-The gain is therefore distributed across the full label space rather than concentrated in a single easy row.
+Under linear probing, shown in Figure 4(a), the most pronounced classwise advantage appears on RGZ DR1: the selected model strengthens all six diagonal entries, reaching 92%, 81%, 78%, 85%, 72%, and 87% from 1C1P through 3C3P.
+The gain is therefore distributed across the full label space rather than concentrated in a single row.
 MiraBest is also favorable in this comparison, with the selected model reaching 68% on FR I and 72% on FR II, ahead of both ViT-MAE and DINOv2 (Registers).
@@
 On MiraBest, all three models are tied at 70% on FR I, while the selected model matches the strongest FR II row at 75%.
-LoTSS DR2 remains the clearest limitation: the selected model is strongest only on FR II (88%) and trails at least one baseline on the remaining rows.
+LoTSS DR2 remains the most evident limitation: the selected model is strongest only on FR II (88%).
 RGZ DR1 remains partly favorable, but the gain is more localized than under probing: the selected model matches the best 1C1P row at 93% and leads on 1C2P and 2C2P (90% and 94%), whereas DINOv2 (Registers) remains stronger on 1C3P, 2C3P, and 3C3P.
@@
 The branch comparison indicates that additional stages are not uniformly beneficial across settings.
-Reconstruction-only continued pretraining is the weakest of the three branches and often underperforms the ViT-MAE starting point.
+Reconstruction-only continued pretraining is the weakest branch and often underperforms the ViT-MAE starting point.
 Masked reconstruction alone is therefore not sufficient to improve transfer in this setting.
-Introducing contrastive learning changes the picture substantially: contrastive-only training already produces large gains over phase 1 alone, especially under linear probing.
+Introducing contrastive learning alters the comparison substantially: contrastive-only training already produces large gains over phase 1 alone, especially under linear probing.
 Across the branch comparison, the two-stage branch is the most consistently top-performing STRADAViT variant, even though a small number of settings—most notably fine-tuned RGZ DR1—are matched or marginally exceeded by contrastive-only training.
-The comparison against the actual initialization is informative in two respects.
-First, relative to ViT-MAE, the strongest two-stage models improve on all linear-probe settings and on fine-tuned MiraBest and RGZ DR1.
-Second, the branch comparison shows that most of this benefit comes from the contrastive objective, while the reconstruction stage contributes a smaller, dataset-dependent benefit.
-Phase 1 is best interpreted as a domain-adaptive warm start or regularizer for the subsequent contrastive stage, not as the main source of representation quality by itself.
+Relative to ViT-MAE, the strongest two-stage models improve on all linear-probe settings and on fine-tuned MiraBest and RGZ DR1.
+Most of this benefit comes from the contrastive objective, while the reconstruction stage contributes a smaller, dataset-dependent benefit and functions primarily as a domain-adaptive warm start for the subsequent contrastive stage.
 Two factors appear most influential: contrastive losses and ROI-aware view generation.
@@
 Register-token effects are also more favorable inside the STRADAViT pipeline than in the off-the-shelf DINO baselines, especially on LoTSS DR2 and RGZ DR1 under probing, although the gain is not uniform across every branch and dataset.
-The DINOv2-Base(R)-initialized ablation is consistent with this pattern: it indicates portability of the contrastive adaptation recipe to a stronger starting point, but the resulting gains remain small and dataset-dependent, with the same LoTSS DR2 limitation that appears across the adapted models more generally.
+The DINOv2 initialization ablation is consistent with this picture: it confirms portability of the contrastive adaptation recipe to a stronger starting point, but under the same recipe the adapted register-based DINOv2 initialization becomes preferable to the non-register initialization, despite the opposite ordering in the off-the-shelf linear-probe baselines.
+The resulting gains remain small and dataset-dependent, with the same LoTSS DR2 limitation that appears across the adapted models more generally.
 The view generator is strongly object-centric.
@@
 LoTSS DR2 and MiraBest contain larger-scale extended structure that benefits from larger-view sampling designed for extended source morphology in radio astronomy.
+This bias is consistent with the weaker and less uniform LoTSS DR2 outcomes, where extended structure is less well matched by the present view generator.
 11 Conclusions and Future Work
 This paper studies how to adapt self-supervised ViT continued pretraining to sparse, heterogeneous radio astronomy imaging.
-The main result is that masked reconstruction alone is insufficient, contrastive learning is the main source of transfer gain, and the two-stage reconstruction-to-contrastive recipe is the most consistent STRADAViT configuration overall.
-The targeted DINOv2 ablation indicates that this adaptation recipe is not specific to the ViT-MAE starting point, but the ViT-MAE-based two-stage checkpoint remains the preferred STRADAViT release because it combines competitive transfer with substantially lower token count and lower downstream cost than the DINOv2-based alternative.
-In that sense, STRADAViT is best interpreted as a domain-adapted starting point for radio astronomy vision backbones: it provides a better practical replacement for out-of-the-box ViTs in frozen-backbone transfer, and remains favorable in most fine-tuning settings without implying uniform superiority in every dataset or class.
+Masked reconstruction alone is insufficient, contrastive learning is the main source of transfer gain, and the two-stage reconstruction-to-contrastive recipe is the most consistent STRADAViT configuration overall.
+Evidence is strongest under frozen-backbone evaluation, while the fine-tuning results remain favorable in most settings.
+The targeted DINOv2 ablation shows that the adaptation recipe is not specific to the ViT-MAE starting point, but the ViT-MAE-based two-stage checkpoint remains the selected STRADAViT release because it combines competitive transfer with substantially lower token count and downstream cost than the adapted DINOv2 alternatives.
+STRADAViT therefore serves as a domain-adapted starting point for radio astronomy vision backbones rather than as a uniformly dominant model on every dataset or class.
 We release this ViT-MAE-based checkpoint on Hugging Face [^1] https://huggingface.co/ISSA-ML/stradavit-base.
@@
 Future work should test stronger teacher–student pretraining schemes, especially DINO-style variants built on anchored radio astronomy-aware view generation, together with larger-view regimes that preserve extended emission more consistently in pretraining.
-Further iterations of the on-the-fly data-curation and view-generation pipeline are also warranted, especially through augmentation and sampling schemes that better match the heterogeneity of the downstream datasets rather than remaining predominantly RGZ-like, while maintaining the general recipe presented here.
-Such refinements should improve the cross-dataset balance of the learned representations and are therefore a plausible route to more uniform downstream gains.
-The strong linear-probe gains also motivate evaluation in fixed-backbone settings, including detection and segmentation pipelines in which STRADAViT can serve as a domain-adapted feature extractor in place of weaker non-specialized ViT backbones.
-In parallel, the stable downstream gains observed under full fine-tuning justify further evaluation of task-specific models built directly on top of STRADAViT.
+Further iterations of the on-the-fly data-curation and view-generation pipeline are also warranted through augmentation and sampling schemes that better match downstream heterogeneity rather than remaining predominantly RGZ-like.
+The strong linear-probe gains motivate evaluation in fixed-backbone settings, including detection and segmentation pipelines, while the stable downstream gains under full fine-tuning justify evaluation of task-specific models built directly on top of STRADAViT.
 Additional priorities are broader cross-survey robustness tests, explicit out-of-domain transfer, and repetition of the same controlled pipeline from stronger off-the-shelf initializations.
 Acknowledgements
-The authors acknowledge Xjenza Malta for funding this study under the Research Excellence Programme of 2024 [REP-2024-19].
+The authors acknowledge Xjenza Malta for funding this study under the STRADA project through the Research Excellence Programme of 2024 [REP-2024-19].
 The authors thank Alessio Magro and Kristian Grixti, Institute of Space Sciences and Astronomy, University of Malta, for technical support and for maintaining the GPU computing infrastructure used for the training and evaluation runs in this study.